### v1

# ImageNot: A contrast with ImageNet preserves model rankings

Abstract
We introduce ImageNot, a dataset designed to match the scale of ImageNet while differing drastically in other aspects. We show that key model architectures developed for ImageNet over the years rank identically when trained and evaluated on ImageNot to how they rank on ImageNet. This is true when training models from scratch or fine-tuning them. Moreover, the relative improvements of each model over earlier models strongly correlate in both datasets. We further give evidence that ImageNot has a similar utility as ImageNet for transfer learning purposes. Our work demonstrates a surprising degree of external validity in the relative performance of image classification models. This stands in contrast with absolute accuracy numbers that typically drop sharply even under small changes to a dataset.
1 Introduction
The Achilles heel of machine learning is its apparent lack of robustness. Models developed in one domain consistently perform worse in new domains. A slew of important—but essentially negative—empirical results demonstrated that even small changes to a dataset can impact the accuracy of machine learning models dramatically (Recht et al., 2019; Taori et al., 2020; Miller et al., 2021). There is good reason to fear that machine learning lacks external validity (Torralba and Efros, 2011; Liao et al., 2021).
Against this pessimistic backdrop, we show that key qualities of machine learning models do generalize to a surprising degree. The starting point of our work is a thought experiment. How different a dataset from ImageNet would have supported essentially the same model developments over the years? ImageNet was at the center of the deep learning revolution in computer vision of the 2010s. Numerous deep architectures were developed specifically with ImageNet in mind. It’s, therefore, reasonable to conjecture that a different benchmark would have spurred different model developments. Our investigation suggests otherwise.
To argue this point, we create a contrast with ImageNet, a new dataset called ImageNot. ImageNot has 1000 classes with 1000 examples per class, just like ImageNet’s ILSVRC-2012 distribution, but ImageNot differs in key characteristics associated with ILSCRC-2012 (Russakovsky et al., 2015; Deng et al., 2009). No human annotators provided any labels. ImageNot pictures instead come from noisy web-crawled data selected based on the image captions and class names alone. While the classes in ILSVRC-2012 have strong concept overlap, the ones we chose for ImageNot are deliberately arbitrary and unrelated. Given the concerns about model architectures overfitting to ImageNet, one expects a new dataset like ImageNot designed to contrast ImageNet to give a range of results that diverge from those associated with ImageNet. Surprisingly, ImageNot turns out to share many of ImageNet’s qualities in terms of model training, comparison and ranking, fine-tuning, and transfer learning.
1.1 Our contributions
(a) Model rankings with models initialized with random weights, trained on a dataset and evaluated on a held-out test set. We note that on CIFAR10, the error bars for EfficientNet-V2-L (Tan and Le, 2019) and ConvNext-L (Liu et al., 2022) overlap, so the rankings may still be exactly preserved.
(b) Relative progress of each model architecture. For a set of test accuracies $X$, relative progress is $X_{i}/X_{\text{AlexNet}}$. This figure is reproduced in Figure 8(b).
Figure 1: Final Model Ranking and Relative Improvement.
Our primary contribution is an in-depth investigation into the external validity, robustness, and reproducibility of model development progress in machine learning. Our work spans the creation of a new dataset and the compute-intensive training and evaluation of numerous architectures on the new dataset.
Creating ImageNot.
We create the ImageNot dataset from the LAION-5B dataset, a public collection of 5 billion image-text pairs crawled from the web and filtered by the CLIP model (Schuhmann et al., 2022; Radford et al., 2021). Although CLIP is a model trained with knowledge of ImageNet, the sheer size of LAION-5B ensures that images in LAION-5B are not already too similar to ImageNet images. We choose 1000 classes from the WordNet lexical database while avoiding nouns close to those included in ImageNet. For each class, we pick 1000 images from LAION-5b based on caption-class similarity alone. We use Roberta text embeddings for this purpose. As a text-only model, Roberta was never trained on ImageNet. This mitigates the concern that our selection of images from LAION-5B might be biased in favor of ImageNet-like instances. We give a detailed analysis of the differences between ImageNet and ImageNot.
ImageNot preserves model rankings and relative improvements.
Our primary empirical finding is that ImageNot has the same image classification model rankings as ImageNet. This finding is important because it provides evidence that the model architecture improvements centered around ImageNet could have occurred similarly for a different dataset than ImageNet. We consider 6 representative model architectures from AlexNet to ConvNeXt-L (Krizhevsky et al., 2012; Liu et al., 2022). We train from scratch and evaluate each model on ImageNot. We observe that the final ranking is the same as on ImageNet. Moreover, the improvement that each model makes relative to prior models is roughly the same on ImageNot as it is on ImageNet. Figure 1 summarizes these results. The remarkable consistency of relative improvements suggests an intriguing counterfactual. Had ImageNot been the benchmark at the time of a new image classification model development, it could have induced the same improvements as ImageNet.
A major hurdle in executing this comparison is that architectures developed for ImageNet do not readily train well on ImageNot. Adapting each architecture to ImageNot requires some amount of hyperparameter tuning. To level the playing field between different architectures, we set careful a priori guardrails around hyperparameter tuning. Specifically, we only tune a small set of hyperparameters that is the same for all models. Moreover, all models get an equal tuning and training budget. These guardrails avoid confirmation bias and limit the researcher’s degrees of freedom.
ImageNot pretraining and transfer learning.
ImageNet has proved useful for pretraining and transfer learning (Kornblith et al., 2019). We wanted to know if the same is true for ImageNot. Toward this end, we conducted two experiments for each model:
  1. 1.
    Pretrain on ImageN[o/e]t, finetune all layers on CIFAR-10, evaluate on CIFAR-10.
  2. 2.
    Pretrain on ImageN[o/e]t, finetune last layer on CIFAR-10, evaluate on CIFAR-10.
We find that a model’s test accuracy on ImageN[o/e]t is a good predictor of fine-tunability and transferability. These experiments suggest that ImageNot shares with ImageNet its utility for pretraining and transfer learning.
To summarize, rankings of image classification models are surprisingly stable across significantly different datasets. Prior work had observed that small variations in a data-generating process preserve model rankings. Our work adds a more extreme contrast. Even significant deliberate deviations from a dataset can maintain essential characteristics of model training and comparison. In particular, our findings suggest that the external validity of machine learning is, in important ways, significantly greater than previously assumed.
Code.
2 Discussion of related work
The external validity of benchmark results for image classification has long been of concern. Indeed, benchmarks often contain biases that do not generalize to the data observed in the real world. Ponce et al. (2006) provide evidence of the narrow range of variation within existing benchmarks in the early stages of image classification research. Subsequently, Torralba and Efros (2011) compare multiple well-known image classification datasets, including SUN09 (Xiao et al., 2010), PASCAL Visual Object Classes (VOC) (Everingham et al., 2010), and Caltech-101 (Fei-Fei et al., 2004). They find evidence of selection bias, capture bias, and negative set bias in these datasets.
The desire for a more representative benchmark for real-world image classification tasks gave birth to ImageNet, a large-scale image database (Deng et al., 2009). The extensive scale and relatively meticulous curation of ImageNet were thought to alleviate some of the limitations identified in earlier datasets, launching over a decade of active development in image classification, largely centered around improving accuracy on the ImageNet dataset. Still, there are concerns about the external validity of ImageNet results. Tsipras et al. (2020) studied the design choices of ImageNet and their effect on alignment between the ImageNet dataset and the broader underlying object recognition task. They concluded that ImageNet contains some inherent biases that cause misalignment. Despite this misalignment, they argue, improvements in multiclass classification on ImageNet still translate to the underlying object recognition task. However, they raised the issue that these biases make it difficult to identify actual improvements in object recognition. Xiao et al. (2020) identify one of these biases in that models tend to rely on the background of ImageNet images, which are correlated with the corresponding labels.
Recht et al. (2019) investigated whether ImageNet classifiers generalize to a newly curated ImageNet test set—recall that by ImageNet, we generally refer to the ILSVRC 2012 ImageNet challenge. They found that accuracies decrease strongly from the original the new test set, in the range of 11%-14%. Despite these findings, Recht et al. (2019)’s results showed that model rankings did not significantly change between the two ImageNet test sets, meaning, if model a performed better than model b on the original ImageNet test set, it also performed better on the new ImageNet test set. This speaks to the stability of model rankings under a slight distribution shift (new versus old test set) but leaves the question of how stable model rankings are under more drastic changes to the dataset. Kornblith et al. (2019) showed that for a set of popular image classification architectures, improvements on ImageNet are often a good predictor of transfer learning performance on other image classification datasets. In addition, beyond image classification, Huang et al. (2017b) showed that improvements in image classification on ImageNet often also predicted improvements in other computer vision tasks, such as image segmentation.
While considerable research on the generalizability of ImageNet models has been centered around changes in accuracy under distribution shift, this work proposes a different perspective. We argue that for benchmarking purposes, providing robust model rankings is a more critical property of an effective benchmark than preserving accuracy or similar metrics. This, however, does not discount the importance of metrics such as accuracy on a well-thought-out and well-curated dataset.
In Mark Liberman’s seminal 2015 lecture on reproducible research and the common task method, he described the role of benchmarks as an aligning mechanism for measuring progress (Liberman, 2015). We distinguish between two related functions of benchmarks in measuring progress: (i) Ranking, i.e., measuring relative progress, (ii) quantifying, i.e., absolute metrics such as accuracy in image classification. Although metrics can be useful, we argue they can be problematic targets for gauging external validity. Consider two identical benchmarks where we inserted a different fraction of random labels in each benchmark. The accuracy numbers may, therefore, be completely different, while the two benchmarks fundamentally measure the same thing. We, therefore, focus on ranking as the primary gauge of external validity.
The observed progress on image classification can be well summarized by a set of models that span the active development of model architectures around ImageNet: AlexNet (Krizhevsky et al., 2012), VGG (Simonyan and Zisserman, 2014), ResNet (He et al., 2016), DenseNet (Huang et al., 2017a), EfficientNet (Tan and Le, 2019), and ConvNeXt (Liu et al., 2022). While other models were also developed during the so-called ImageNet era, we focus on models that were most impactful and represented a notable shift in architecture design – we include a discussion in Appendix A. Specifically, we considered the availability of model implementations across hubs such as TorchHub (PyTorch Team, 2024), TensorFlow (Abadi et al., 2015), and Huggingface (Wolf et al., 2020) to be an indicator of popularity and, therefore, impact. We note that while newer architectures such as Vision Transformers  (Devlin et al., 2019; Dosovitskiy et al., 2020) are now commonplace in the literature, we only consider architectures developed to improve ImageNet accuracy in our study.
Many variants of ImageNet have been created to test the robustness of ImageNet models. Hendrycks and Dietterich (2019) introduced ImageNet-C/P, ImageNet with common corruptions.  Hendrycks et al. (2020) introduced ImageNet-R, stylized distribution shifts on ImageNet. Li et al. (2022) introduced ImageNet-W, ImageNet with watermarks, a learning shortcut. Shirali and Hardt (2023) introduced LAIONet, an intended recreation of ImageNet from the LAION-400M dataset, in order to identify key differences between ImageNet and LAIONet.
Our aim is distinctly different from these works. Rather than studying variants of ImageNet, we create an entirely different dataset. We aim to evaluate if this completely different dataset suggests the same model development trajectory, i.e., incremental rankings, as ImageNet. Accordingly, we introduce a new dataset, ImageNot, with entirely different classes and a unique curation process, drawing inspiration from LAIONet’s data-generating process (Shirali and Hardt, 2023). A distinct difference between our work and these previous works is that we train and test on this new dataset. This contrasts to (Recht et al., 2019; Shirali and Hardt, 2023), who compare test accuracies of the same ImageNet trained models on two different test sets with ImageNet classes. Our results compare the test accuracy of ImageNet-trained models on ImageNet’s test set with the test accuracy of ImageNot-trained models on ImageNot. This removes the confound of ImageNet training and the ImageNet classes, better isolating the external validity of model architectures independent of their training data. Additionally, we developed ImageNot rather than other datasets to ensure the strength of its contrast with ImageNet.
3 Constructing ImageNot
Figure 2: ImageNot caption x synset-gloss RoBERTa embeddings similarity for each class. We show the minimum, mean, median, and maximum similarity distribution across the 1000 classes.
Figure 3: ImageNot caption x synset-gloss RoBERTa embeddings similarity distribution for each class. Each distribution corresponds to similarities for each class.
We develop ImageNot, a dataset of the same scale as ImageNet’s ILSVRC-2012 – 1000 classes with 1000 examples per class – but without key design choices that have been thought critical in ImageNet’s effectiveness (Russakovsky et al., 2015). First, ImageNot has deliberately arbitrary classes constrained only to be strictly different from those in ImageNet. Second, we do not utilize any human annotators. There were no humans in the loop in selecting specific image and label pairs. We create the ImageNot dataset from the LAION-5B dataset, an academic, public massive scale collection of 5 billion image-text pairs crawled from the web and filtered by the CLIP model (Schuhmann et al., 2022; Radford et al., 2021). This is in contrast with the ILSVRC 2012 challenge, a subset of the ImageNet database which was constructed from Flickr image search results (Deng et al., 2009).
Similar to ImageNet, ImageNot classes correspond to nouns in WordNet synsets found in LAION-5B captions, where WordNet is a large lexical database of English (Fellbaum, 1998; Miller, 1995; Hardeniya et al., 2016). The following definitions are important for describing WordNet. A given word has a set of synsets, where a synset is a sense of a word and has a corresponding short definition called a gloss. For example, ‘car’ has 2 synsets with corresponding glosses: (i) “a motor vehicle with four wheels; usually propelled by an internal combustion engine” and (ii) “a wheeled vehicle adapted to the rails of a railroad.” Additionally, each synset consists of a set of synonyms called lemmas. The lemmas of ‘car’-(i) include automobile, motorcar, while the lemmas of ‘car’-(ii) include railcar, railwaycar. A WordNet tree represents a semantic hierarchy of synsets via an ‘IS-A’ structure, i.e., a synset is a ‘type’ of its synset parent, e.g., mammal $\rightarrow$ placental $\rightarrow$ carnivore $\rightarrow$ dog $\rightarrow$ working dog $\rightarrow$ husky.
Each class in ImageNot represents a unique noun synset. At a high level, for each class, we select images from LAION-5B whose captions are similar to the synset corresponding to the class. We utilize RoBERTa (Liu et al., 2019) embeddings, an optimized version of BERT (Devlin et al., 2019), to estimate similarity between text sequences. Since RoBERTa was never trained on ImageNet, this selection mechanism mitigates the concern that our image selection process may be biased in favor of ImageNet-like images.
We note that creators of LAION-5B’s only included image and caption pairs that are sufficiently similar based on CLIP embeddings (Contrastive Language–Image Pre-training (Radford et al., 2021)), a multimodal deep learning model (Schuhmann et al., 2022).
Constructing ImageNot.
To choose the 1000 synsets (classes) that are included in ImageNot, we do the following (employing similar mechanics as Shirali and Hardt (2023)):
  1. 1.
    Find all noun synsets in the LAION-5B captions.
  2. 2.
    Remove all synsets that belong to any ImageNet synset/lemmas.
  3. 3.
    Filter synsets by size, dropping any candidate synset with too few image-caption pairs.
  4. 4.
    Find the minimal subtree of the WordNet synset tree that contains all of the remaining candidate synsets and select only the synsets that are leaves of this subtree.
  5. 5.
    Filter out synsets considered ‘Not Safe for Work’ (NSFW) by Song et al. (2021)’s NSFW classifier.
  6. 6.
    Find the subset of captions such that no caption contains multiple ImageNot synset’s lemmas
  7. 7.
    Filter out synsets with an average top n caption x synset-gloss pair RoBERTa embeddings similarity (normalized inner product) less than 0.82  (Liu et al., 2019). We found 0.82 to be the highest reasonable choice to allow for sufficient examples in enough classes (Figure 3).
  8. 8.
    Finally, select the top 1000 classes ordered by the average similarity between caption x synset-gloss pairs (Liu et al., 2019) (desideratum 4).
Finally, we obtain ImageNot, a large-scale dataset that contrasts ImageNet. ImageNot has an average caption-label definition similarity of 0.91. Next, we compare ImageNot to ImageNet. ImageNot classes can be found at https://github.com/olawalesalaudeen/imagenot.
3.1 ImageNot vs. ImageNet
Figure 4: Synset depth in WordNet tree.
Figure 5: Class Diversity. Average intra-class image similarity – inner product between normalized CLIP embeddings.
Figure 6: Image Variance. Average inter-class image similarity – inner product between normalized CLIP embeddings.
In this section, we delve into the differences between ImageNot and ImageNet, focusing on their design aspects. The primary distinctions lie in two areas: (i) the classes in ImageNot are arbitrarily different from those in ImageNet, and (ii) the attribution of a synset to an image in ImageNot is conducted without human intervention. We begin by examining the key differences between the two datasets.
The ‘Not’ in ImageNot.
The first significant not is the hierarchy of synsets, based on the WordNet tree. ImageNet is constructed with synsets representing a densely populated subset of the WordNet tree. For instance, ImageNet contains 147 categories of dogs. In contrast, ImageNot is not designed to have densely populated subsets of the WordNet tree. Moreover, ImageNot synsets tend to be higher up in the WordNet tree (Figure 4) with ImageNot synsets having a median depth of 8 while ImageNet synsets have a median depth of 10 (more details about the WordNet subtrees can be found in Appendix B.1 Table 3). One consequence of this difference in median depth is that the average diversity of images in an ImageNot synset is higher than that of ImageNet. We illustrate the distribution of intra-class similarity within each dataset in Figure 6, where ImageNet exhibits higher intra-class diversity than ImageNot.
The second significant not is the accuracy of image-synset pairs. In ImageNet, an average precision of 99.7% is achieved on a total of 80 synsets randomly sampled at different depths of the WordNet tree. Candidate images were collected from the internet by querying image search engines, and then human annotators were employed to verify each candidate image for a given synset, yielding relatively accurate image-synset pairs. In contrast, we deliberately make no such cleaning steps in ImageNot. We rely entirely on similarities between embeddings from CLIP (image and caption similarities) and RoBERTa (caption and class definition similarities) to select class and image pairs. Consequently, ImageNot’s labels are far more noisy than ImageNet, which utilizes human annotators. This noisiness is noticeable in the accuracy observed when training on ImageNot – the best model achieves around 50%, which is drastically better than chance (0.1%), but lower than what the best model achieves on ImageNet (around 85%). We also see the effect of this process in the distribution of inter-class image similarities between ImageNot and ImageNet illustrated in Figure 6; the similarity in images corresponding to an ImageNet class tends to be higher than those in an ImageNot class.
Figure 7: Number of synset siblings.
Similarities in ImageNot and ImageNet WordNet subtrees.
Despite the differences in ImageNot, the subset of the WordNet trees spanned by their synsets are structurally similar. They share approximately the same number of nodes and edges, have similar depth, density and modularity, and share a 23% overlap in the parents of the datasets’ synsets – illustrated in Figure 7 (more details about the WordNet subtrees can be found in Appendix B.1 Table 3).
4 ImageNot preserves model rankings
(a) Models initialized with random weights. Train models on a dataset. Evaluate models on the same dataset’s held-out test set.
(b) Relative progress of each model architecture. For a set of test accuracies $X$, relative progress is $X_{i}/X_{\text{AlexNet}}$.
(c) Test accuracy compared to ImageNet test accuracy for each model architecture.
Figure 8: ImageNet vs. ImageNot. We also show 95% confidence intervals for the regression estimates. Robust corresponds to average accuracy across many test-time distribution shifts.
Our experimental design is centered around evaluating whether ImageNot maintains the model rankings established by ImageNet. Specifically, we want to know: if a new architectural advancement was evaluated on ImageNot instead of ImageNet, would it still be recognized as a significant development? Our focus is on a collection of model architectures that represent significant milestones in the evolution of ImageNet models. This includes AlexNet (Krizhevsky et al., 2012), VGG (Simonyan and Zisserman, 2014), ResNet (He et al., 2016), DenseNet (Huang et al., 2017a), EfficientNet (Tan and Le, 2019), and ConvNeXt (Liu et al., 2022). We motivate this set in section 2. For each model, we select the configuration that achieved the highest reported performance on ImageNet, such as ResNet152 over ResNet18. These models are then trained and evaluated on ImageNot, allowing us to directly compare their rankings with those obtained on ImageNet.
To test each model architecture, we train a version on ImageNet and another on ImageNot and evaluate them on the corresponding test set. We sort the test accuracy to establish a definitive order of performance in object recognition tasks, which we refer to as the ‘model ranking.’
We also investigate model rankings on the CIFAR10 dataset, which consists of 60,000 32x32 colored images of 10 classes, with 5000 train images per class and 1000 test images per class (Krizhevsky et al., ; Krizhevsky et al., 2009). The learning task is 10-class image classification – Airplane, Automobile, Bird, Cat, Deer, Dog, Frog, Horse, Ship, and Truck. We note that ImageNet contains 50 test images per class (Deng et al., 2009) while ImageNot contains 100.
We consider three types of rankings: (i) Standard. We initialize each model with random weight, then train each model on a dataset, and finally, we evaluate each model on the same dataset’s test set. (ii) Fine-Tuning. We initialize each model with pre-trained on ImageN[o/e]t weights, then we fine-tune the model end-to-end (all weights updated) on CIFAR10, and finally, we evaluate it on CIFAR10’s test set. (iii) Transfer Learning. We initialize each model with pre-trained ImageN[o/e]t weights, then we update only the model’s last fully connected layer while training on CIFAR10, and finally, we evaluate it on CIFAR10’s test set.
4.1 Model development
ImageNet. Our reference rankings are derived from pre-trained models available on Torch Hub (Marcel and Rodriguez, 2010; Paszke et al., 2019). These models establish a ranking, which we consider our factual observations. In decreasing order of performance, the ranking is as follows: EfficientNet V2 Large, ConvNeXt Large, ResNet 152, DenseNet 161, VGG 19, ResNet 18, and AlexNet. For additional specifics regarding these pre-trained models, please refer to Appendix A.1.
Training.
We also utilize the implementation and pre-trained model weights from torch hub (Marcel and Rodriguez, 2010; Paszke et al., 2019), when applicable. For the first ranking category (no pretraining), we initialize each model with randomized weights and train end-to-end, meaning all model weights are updated with each backward pass.
On avoiding self-fulfilling prophecies.
There is a subtle pitfall that one may fall into if not careful – the inadvertent confounding effect of knowing the desired ImageNet ranking. Specifically, an experimenter aware of the expected ranking might unconsciously adjust the process until a particular model is in the right place. This iterative tuning could falsely lead to a preservation of the known rankings. To mitigate this risk, we implemented several a priori restrictions on the experimenter’s approach for each ranking modality:
  1. 1.
    The experimenter can only tune the learning rate, learning rate scheduler hyper-parameters, and weight decay constant.
  2. 2.
    There is a shared hyper-parameter set across all models.
  3. 3.
    There is a fixed # of training runs that can be launched for the same maximum epoch.
  4. 4.
    The fixed hyper-parameter set can only be changed (expanded) if some model(s) do not train.
See further discussion in Appendix A.2. This cautious approach introduces the primary limitation of our work. To avoid over-tuning and maintain objectivity, we refrain from employing extensive model tuning or leveraging all modern performance-enhancing techniques, such as data augmentation, label smoothing, etc. As a result, it is possible that we do not end up with the optimal training configuration for each model (Figure 8(c)).
4.2 Model Rankings
ImageNot preserves model rankings and relative improvements.
Figure 8(a) illustrates that the model rankings between ImageNot and ImageNet are remarkably consistent. In Figure 8(b), we compare each model’s performance improvement relative to AlexNet, the model that kick-started the deep learning era on ImageNet in 2012. This comparison reveals a notable parallel in the progression of accuracy improvements across both ImageNet and ImageNot. Not only do the ordinal rankings of models remain unchanged, the degree of performance enhancement is also consistent. In simpler terms, this suggests that if a new architecture like EfficientNet (Tan and Le, 2019) had been introduced in a different universe focused on ImageNot instead of ImageNet, we would have witnessed a comparable impressive advancement. The external validity of ImageNet improvements holds! The ImageNet vs. ImageNot test accuracies have a linear fit with a slope of 0.46 and an adjusted $R^{2}$ of 0.98. ImageNet vs. CIFAR10 test accuracies have a linear fit with a slope of 0.18 and an adjusted $R^{2}$ of 0.93. We note that these slopes should not be interpreted the same way as those in Recht et al. (2019) because we compare test accuracies across datasets, i.e., different ranges in accuracy across models.
(a) ImageN[o/e] pre-trained models. Train on CIFAR10. Evaluate on the CIFAR10. Similar to previous rankings, changes occur between architectures with overlapping error bars.
(b) Fine-tuned on CIFAR10 test accuracy compared to standard ImageNet test accuracy for each model architecture.
(c) Transfer Learning (fine-tune only last fully-connected layers) on CIFAR10 test accuracy compared to standard ImageNet test accuracy for each model architecture.
Figure 9: ImageNet vs. ImageNot. Fine-tuning and Transfer Learning. Robust corresponds to average accuracy across many test-time distribution shifts. We also show 95% confidence intervals for the regression estimates.
Figure 8(c) highlights the limitations of relying solely on the external validity of accuracy as the primary metric for the effectiveness of a benchmark. One may look at the drop in accuracy from ImageNet to ImageNot and conclude that ImageNet results have no external validity but this would be a rather narrow and inaccurate conclusion. From the standpoint of benchmarking, the more relevant observation and notion of external validity is that ImageNet accuracy is a strong predictor of ImageNot accuracy (Figure 8(c)), despite the significant difference between the two datasets.
For each CIFAR10 ranking, we do three random restart restarts and compute the standard errors for each model’s test accuracy. Given that we opted to include architectures that achieve very similar accuracies on the ImageNet test set (such as EfficientNet-V2-L/ConvNeXt-L or ResNet152/DenseNet161), we expect potential training variance to affect the model, particularly given the size of CIFAR10. For instance, when training CIFAR10 from randomly initialized weights, EfficientNet-V2-L shows a slightly lower average test accuracy than ConvNeXt-L across the random restarts. However, their error bounds intersect, so the ranking is preserved up error margins.
Our methodology restricts experimenters such that once a hyperparameter search space and the number of samples from the search space are fixed, they cannot be adjusted unless a model does not train. This approach safeguards the rankings from confirmation bias. However, we note that it also has some notable limitations, e.g., models with similar capabilities with overlapping error bars.
ImageNot improvements yield improvements in fine-tuning and transfer learning.
When fine-tuning pre-trained ImageN[o/e]t models on CIFAR10, we observe that model rankings are still preserved (Figure 9(a)). Within a fixed hyperparameter set, however, we find that some of the models, particularly those with the highest accuracies, have high variance. Nevertheless, both the upper bounds and lower bounds of the error margins increase as the accuracy on ImageNet increases (Figure 9(b)).
Furthermore, despite the arbitrary nature of the ImageNot dataset, we observe that the representations it induces are still transferable – the accuracy of a model architecture on ImageNot is still a good predictor of the accuracy achieved by retraining the last layer of a model pre-trained on ImageNot on CIFAR10, (Figure 9(c)). However, we note that the ImageNot features are notably more noisy than ImageNet features when fine-tuning. This suggests that the noisiness of ImageNot does have some negative downstream effects on transfer learning. Nevertheless, we generally observe a positive trend in ImageNot test accuracy and transferability to CIFAR10.
Rankings are preserved under robustness.
We also find that model rankings remain consistent when we look at average performance across 17 distribution shifts applied to the ImageN[o/e]t test set, e.g., random blurring, cropping, and rotation. In this setting, our findings remain the same on average across these distribution shifts but with a consistent decrease in test accuracy for each model. This further shows that metrics can be a misleading definition of external validity as we observe changes in accuracy across different shifts. More details on rankings under robustness are in Appendix B.2.
Safety and ethical limitations of web-crawled datasets.
While our results suggest that web-crawled datasets with minimal human intervention may be sufficient for benchmarking progress in architecture development, other ethical considerations are a concern. As illustrated by Thiel (2023), the internet contains many illegal images and images of illegal activity. Particularly, Thiel (2023) find evidence of Child Sexual Abuse Material (CSAM) in LAION-5B. We guard against this by using a NSFW language classifier (Song et al., 2021) to remove synsets that may be associated with unethical images on top of the default LAION-5B NSFW filter. We did not manually inspect every image in ImageNot, but we employed our model-based filtering and design choices to avoid problematic images as best as possible. Nevertheless, we do not release the images or URLs for ImageNot, but we provide ImageNot class names, a detailed description of our generation procedure, and our code so that our results can be reproduced.
5 Conclusion
Despite the significant differences in ImageNet and ImageNot, we find consistent preservation of model rankings across both datasets. Moreover, this consistency extends to the fine-tunability of pre-trained models and the transferability of pre-trained image embeddings. For benchmarking, we propose that providing robust model rankings should be a hallmark of a good benchmark. Furthermore, our results suggest that concerns about the external validity of model developments on ImageNet may be overstated.
Many domain experts attribute the rapid advancement in object recognition models to the ImageNet benchmark. From this work, two takeaways emerge to guide and motivate future benchmarks. First, a benchmark that produces robust model rankings, even if accuracy degrades under distribution shifts, should be considered effective. Second, the emphasis on clean labels and accuracy numbers may not be necessary to measure progress effectively, given that the noise is unbiased.
The exact extent to which a benchmark needs to align with real-world image distributions to measure progress effectively remains a question for future work. Our findings indicate that this alignment might need to be less stringent than currently perceived in the machine learning community.
Additional future work includes changing additional design choices of ImageN[o/e]t, e.g., moving away from wordnet noun synsets and using other attributes of or relationships between concepts in images. Another direction for future work is to examine the properties considered in this work for fine-tuning ImageNet and ImageNot-type models on other downstream tasks such as image segmentation.
Acknowledgements
We are indebted to Ali Shirali for providing his code for generating datasets from LAION, which we modified to create ImageNot. We thank André Cruz, Ricardo Dominguez-Olmedo, Florian Dorner, and Katherine Tsai for many helpful discussions on this project and invaluable feedback on an earlier version of this paper. We also thank Sanmi Koyejo for helpful discussions on this project.
Appendix A Architectures and Training
Architecture Development Trajectory.
AlexNet (Krizhevsky et al., 2012) demonstrated the efficacy of deep learning and many architectural advances such as convolutional filters, rectified linear units (ReLUs) as activation functions, and more. Additionally, its relative success on the ILSVRC 2012 is considered to have revitalized interest in neural networks, sparking a wave of research in deep learning. VGG (Simonyan and Zisserman, 2014) then implemented a simpler and deeper architecture with smaller convolutional filters, along with regularization techniques such as dropout, which yielded improved accuracy and generalization. ResNet (He et al., 2016) further improved upon the architecture of these deep learning models with the introduction of residual blocks, which allowed for the training of deeper networks while mitigating the vanishing gradient problem and achieving superior performance to its predecessor. DenseNet (Huang et al., 2017a) proposed to connect each layer to every other layer to maximize information flow between layers. EfficientNet (Tan and Le, 2019) then improved upon its predecessors with compound and systematic scaling of depth, width, and resolution, as well as combinations of novel optimization and learning strategies, without increasing computational cost. ConvNeXt (Liu et al., 2022) aimed to compete with the state-of-the-art performance of Vision Transformers (ViTs; (Dosovitskiy et al., 2020)) by combining components of preceeding ConvNet architectures, e.g., residual blocks and kernel sizes.
A.1 Torch Hub Pretrained Models
We use pretrained models from Torch Hub; we summarize their training details below (see PyTorch Team (2024) for additional details).
Table 1: Default Hyperparameters
Parameter Default Value
batch size 32
epochs 90
lr 0.1
momentum 0.9
weight decay 1e-4
lr step size 30
lr gamma 0.1
AlexNet, VGG19.
Learning rate: 1e-2. Learning rate scheduler: StepLR.
ResNet, DenseNet. Learning rate scheduler: StepLR.
EfficientNet V2 L. Learning rate: 5e-1. Learning rate scheduler: Cosine annealing with linear warmup. Weight decay: 2e-5. EfficientNet V2 L also utilizes data augmentation, label smoothing, and other regularization techniques.
ConvNeXt Large. Learning rate: 1e-3. Learning rate scheduler: Cosine annealing with linear warmup. ConvNeXt also utilizes data augmentation, label smoothing, and other regularization techniques.
For both finetuning and transfer learning from ImageNet, we use these pretrained weights as our starting point.
A.2 Trained Models
We note that for ImageNot, the extent of our parameter tuning is significantly less than these pretrained ImageNet models (Table 2). Despite this, model rankings are still preserved. In general, when we train CIFAR10, we use the same set of hyperparameters, given that we can perform a broader search. This includes training from randomly initialized weights, finetuning, and transfer learning.
For ImageNet, the tuning of these parameters we manual and not search-based – each new hyperparameter was added based on the loss curves of models that did not sufficiently train and domain expertise of which hyperparameters can be changed to allow for training.
Table 2: Hyperparameters search space and Training Budgets for experiment. TL: Transfer Learning; FT: Finetuning; ST: Standard (train/test with randomly initialized weights).
Dataset Hyperparameter Train Style Range Budget
CIFAR10 batch size TL/FT/ST [2048/256/256]* 24
CIFAR10 optimizer TL/FT/ST [SGD, Adam] 24
CIFAR10 lr TL/FT/ST [1e-3, 1e-1] 24
CIFAR10 lr scheduler TL/FT/ST [Cosine Annealing with Warm Restarts] 24
CIFAR10 momentum TL/FT/ST [0.8, 0.99] 24
CIFAR10 optimer warmup steps TL/FT/ST [5, 10, 15, 25, 50] 24
CIFAR10 weight decay TL/FT/ST [1e-6, 5e-3] 24
ImageNot batch size ST [256] 4
ImageNot optimizer ST [SGD] 4
ImageNot lr ST [5e-1, 1e-1, 1e-2, 1e-4] 4
ImageNot lr decay ST [0.1] 4
ImageNot lr scheduler ST [Linear Warm Start + Cosine Annealing] 4
ImageNot momentum ST [0.9] 4
ImageNot optimer warmup steps ST [5] 4
ImageNot weight decay ST [2e-5] 4
(*) For the reported batch size in Table 2, we use [128/32/32] for EfficientNet-V2-L and [256/64/64] for ConvNeXt-large.
We utilize early stopping with a patience of 10 on the basis of held-out validation sample accuracy and set a maximum epoch of 1000 for all training. Additionally, to optimize training efficiency and reduce computational costs, we employ the Asynchronous Hyper-Band Scheduler from Ray Tune (Li et al., 2020; Liaw et al., 2018), which enables us to further early-stop less promising hyperparameters.
Appendix B Additional Results and Discussion
B.1 ImageNet vs. ImageNot WordNet subtree
Table 3: ImageNot and ImageNet (ILSVRC 2012) WordNet tree statistics.
Stat ImageNot ImageNet $\frac{\text{ImageNot}}{\text{ImageNet}}$
Nodes 1877 1860 1.01
Edges 1899 1937 0.98
Min Depth 2 4 0.50
Median Depth 8 10 0.80
Max Depth 18 17 1.06
Density 5.39e-4 5.60-4 0.96
Modularity 0.94 0.93 1.01
Synset Overlap 0% 0%
Parent Overlap 23% 23%
Table 3 summarizes the differences (or lack there of) between properties of the ImageNet and ImageNot WordNet synsets (referenced in section 3.1).
B.2 Robustness
Robustness Ranking.
We employ 17 distinct test-time data augmentations to independently rank model accuracies, all random except for the Identity augmentation. These augmentations are applied to the same set of held-out test examples. They include: (1) Identity, which uses the original image; (2) Gaussian Blur, applying a Gaussian blur to smooth the image and reduce noise; (3) Adjust Sharpness, for modifying image sharpness; (4) Affine, performing affine transformations such as scaling, rotations, and translations; (5) ColorJitter, adjusting brightness, contrast, saturation, and hue; (6) Contrast, altering contrast levels; (7) Equalize, for histogram adjustment and contrast balancing; (8) Grayscale, converting the image to grayscale; (9) Horizontal Flip, flipping the image horizontally; (10) Invert, inverting pixel values; (11) Perspective, applying perspective transformations; (12) Posterize, reducing the bit depth per color channel; (13) Resized Crop, cropping after resizing; (14) Rotation, rotating the image; (15) Sharpness, for sharpness adjustment; (16) Solarize, inverting colors post a threshold; and (17) Vertical Flip, flipping the image vertically.
We take the average across these augmentations to generate our robustness rankings. The model rankings are surprisingly stable on average across these 17 test-time augmentations (Figure 10-12).
Figure 10: Models initialized with random weights. Train on ImageNet. Evaluate on ImageNet test set.
We find that models trained on ImageNet maintain rankings on average when evaluated on the ImageNet test sets with test-time data corruption (Figure 10).
Figure 11: Models initialized with random weights. Train on ImageNot. Evaluate on ImageNot test set.
Additionally, we find that models trained on ImageNot maintain rankings on average when evaluated on the ImageNot test sets with test-time data corruption (Figure 11). However, across the corruptions, the rankings are less stable than what we observed for ImageNet. We hypothesize that this is a consequence of our guardrails on model tuning, which leads to sub-optimal models than those that could have been achieved with more extensive tuning.
Figure 12: Models initialized with random weights. Train on CIFAR10. Evaluate on CIFAR10 test set.
Finally, we find that models trained on CIFAR10 maintain rankings on average when evaluated on the CIFAR10 test sets with test-time data corruption (Figure 12). However, across the corruptions, the rankings are less stable than what we observed for ImageNet. We hypothesize that this is a consequence of our guardrails on model tuning, which leads to sub-optimal models than those that could have been achieved with more extensive tuning.

### v2

@@
 Abstract
-We introduce ImageNot, a dataset designed to match the scale of ImageNet while differing drastically in other aspects.
-We show that key model architectures developed for ImageNet over the years rank identically when trained and evaluated on ImageNot to how they rank on ImageNet.
-This is true when training models from scratch or fine-tuning them.
+We introduce ImageNot, a dataset constructed explicitly to be drastically different than ImageNet while matching its scale.
+ImageNot is designed to test the external validity of deep learning progress on ImageNet.
+We show that key model architectures developed for ImageNet over the years rank identically to how they rank on ImageNet when trained from scratch and evaluated on ImageNot.
 Moreover, the relative improvements of each model over earlier models strongly correlate in both datasets.
-We further give evidence that ImageNot has a similar utility as ImageNet for transfer learning purposes.
-Our work demonstrates a surprising degree of external validity in the relative performance of image classification models.
+Our work demonstrates a surprising degree of external validity in the relative performance of image classification models when trained and evaluated on an entirely different dataset.
 This stands in contrast with absolute accuracy numbers that typically drop sharply even under small changes to a dataset.
@@
 Models developed in one domain consistently perform worse in new domains.
-A slew of important—but essentially negative—empirical results demonstrated that even small changes to a dataset can impact the accuracy of machine learning models dramatically (Recht et al., 2019; Taori et al., 2020; Miller et al., 2021).
-There is good reason to fear that machine learning lacks external validity (Torralba and Efros, 2011; Liao et al., 2021).
+A slew of important—but essentially negative—empirical results demonstrated that even small changes to a dataset can impact the accuracy of machine learning models dramatically (recht2019imagenet; taori2020measuring; miller2021accuracy).
+Consequently, there is good reason to fear that machine learning lacks external validity (torralba2011unbiased; liao2021we).
 Against this pessimistic backdrop, we show that key qualities of machine learning models do generalize to a surprising degree.
@@
 Numerous deep architectures were developed specifically with ImageNet in mind.
-It’s, therefore, reasonable to conjecture that a different benchmark would have spurred different model developments.
+It’s therefore reasonable to conjecture that a different benchmark would have spurred different model developments.
 Our investigation suggests otherwise.
 To argue this point, we create a contrast with ImageNet, a new dataset called ImageNot.
-ImageNot has 1000 classes with 1000 examples per class, just like ImageNet’s ILSVRC-2012 distribution, but ImageNot differs in key characteristics associated with ILSCRC-2012 (Russakovsky et al., 2015; Deng et al., 2009).
+ImageNot has 1000 classes with 1000 examples per class, just like ImageNet’s ILSVRC-2012 distribution, but ImageNot differs in key characteristics associated with ILSVRC-2012 (ILSVRC15; deng2009imagenet).
 No human annotators provided any labels.
@@
 While the classes in ILSVRC-2012 have strong concept overlap, the ones we chose for ImageNot are deliberately arbitrary and unrelated.
-Given the concerns about model architectures overfitting to ImageNet, one expects a new dataset like ImageNot designed to contrast ImageNet to give a range of results that diverge from those associated with ImageNet.
-Surprisingly, ImageNot turns out to share many of ImageNet’s qualities in terms of model training, comparison and ranking, fine-tuning, and transfer learning.
+The central question we ask is: Does ImageNot provide the same model rankings as ImageNet?
+An important difference between our work and previous work is that they investigate the robustness of ImageNet models, models trained on ImageNet.
+In contrast, we investigate the external validity of the core deep learning developments on ImageNet; thus, we remove the ImageNet confound, unlike in previous work.
+The architectures developed on ImageNet are trained from scratch and evaluated on ImageNot.
+More generally, we identify five major confounds that hinder interpretations of previous work in the same context of this work on external validity: Previous work had (i) evaluation sets with same images as ImageNet but with corruptions (Hendrycks2019BenchmarkingNN; hendrycks2021natural; Hendrycks2020TheMF), (ii) the same data sources, Flickr (recht2019imagenet), (iii) the same image classes (recht2019imagenet; barbu2019objectnet; Hendrycks2019BenchmarkingNN; hendrycks2021natural; Hendrycks2020TheMF), and (iv) the same data generative mechanisms, i.e., the procedure of human annotation (recht2019imagenet).
+Furthermore, (v) the same models trained on ImageNet are evaluated (recht2019imagenet; barbu2019objectnet; Hendrycks2019BenchmarkingNN; hendrycks2021natural; Hendrycks2020TheMF).
+While the results of these previous works have been interpreted in the context of external validity of deep learning model development, i.e., scientific progress on ImageNet, these confounds limit the soundness of such interpretations, since they may be responsible for the observed robustness of rankings.
+These confounds leave the question of the external validity of ImageNet as a benchmark for model development progress unanswered.
+Importantly, the goal of this work is not to propose ImageNot as a new general benchmark or replacement for ImageNet.
+ImageNot is a construct, removing these confounds, to address this question and stress test the external validity of model development progress on ImageNet.
+Additionally, our work focuses on benchmarks as a tool for advancing science; there are many other interpretations of benchmarks that one may be interested in, e.g., as instruments for producing valid scores of a model’s capabilities.
+Our demonstration of the robustness of rankings under certain benchmark properties, e.g., noise, should not be assumed to extend to other interpretations of benchmarks.
+In a nutshell, we give a strong affirmative answer to this question of the external validity.
+Key model architectures from the ImageNet era rank identically on ImageNot as they do on ImageNet.
+Moreover, each model’s relative improvement compared with prior models is approximately the same on both datasets.
+Our empirical findings provide strong evidence for the external validity of model rankings.
 1.1 Our contributions
-(a) Model rankings with models initialized with random weights, trained on a dataset and evaluated on a held-out test set.
-We note that on CIFAR10, the error bars for EfficientNet-V2-L (Tan and Le, 2019) and ConvNext-L (Liu et al., 2022) overlap, so the rankings may still be exactly preserved.
-(b) Relative progress of each model architecture. For a set of test accuracies $X$, relative progress is $X_{i}/X_{\text{AlexNet}}$. This figure is reproduced in Figure 8(b).
 Figure 1: Final Model Ranking and Relative Improvement.
+Model rankings and relative improvements hold with models trained from scratch on the respective dataset, and evaluated on a held-out test set.
+For a set of test accuracies $X$, relative progress is $X_{i}/X_{\text{AlexNet}}$.
+ImageNet model accuracies range from 0.57 (AlexNet) to 0.85 (EfficientNet V2 L), while much noisier ImageNot model accuracies approximately range from 0.4 (AlexNet) to 0.6 (EfficientNet V2 L).
+Figure 2: Examples from ImageNet (left) and ImageNot (right).
+ImageNot classes (e.g., ‘cleats,’ ‘batter’) are disjoint and distinct from ImageNet classes (e.g., ‘Irish Terrier,’ ‘Blenheim Spaniel’).
+ImageNot, web-sourced data with automated data selection and labeling, also has greater natural variability and label noise.
 Our primary contribution is an in-depth investigation into the external validity, robustness, and reproducibility of model development progress in machine learning.
-Our work spans the creation of a new dataset and the compute-intensive training and evaluation of numerous architectures on the new dataset.
+We create a new dataset and perform compute-intensive training and evaluation of numerous architectures on the new dataset.
 Creating ImageNot.
-We create the ImageNot dataset from the LAION-5B dataset, a public collection of 5 billion image-text pairs crawled from the web and filtered by the CLIP model (Schuhmann et al., 2022; Radford et al., 2021).
-Although CLIP is a model trained with knowledge of ImageNet, the sheer size of LAION-5B ensures that images in LAION-5B are not already too similar to ImageNet images.
-We choose 1000 classes from the WordNet lexical database while avoiding nouns close to those included in ImageNet.
-For each class, we pick 1000 images from LAION-5b based on caption-class similarity alone.
+We created ImageNot from the LAION-5B dataset, a public collection of 5.8 billion image-text pairs crawled from the web and filtered by the CLIP model (schuhmann2022laion; radford2021learning).
+Although CLIP is trained with knowledge of ImageNet, the sheer size of LAION-5B ensures that images in LAION-5B generally are not similar to ImageNet images.
+We choose 1000 classes from the WordNet lexical database while avoiding nouns close to those in ImageNet.
+For each class, we pick 1000 images from LAION-5B by caption-class similarity alone.
 We use Roberta text embeddings for this purpose.
@@
 We give a detailed analysis of the differences between ImageNet and ImageNot.
+Crucially, our creation of ImageNot avoids confounds in previous work with respect to images and classes, data source, and generative mechanism.
+We provide details instructions for deriving ImageNot from LAION-5B.
 ImageNot preserves model rankings and relative improvements.
@@
 This finding is important because it provides evidence that the model architecture improvements centered around ImageNet could have occurred similarly for a different dataset than ImageNet.
-We consider 6 representative model architectures from AlexNet to ConvNeXt-L (Krizhevsky et al., 2012; Liu et al., 2022).
+We consider 9 representative model architectures from AlexNet to Vision Transformers (krizhevsky2012imagenet; dosovitskiy2020image).
 We train from scratch and evaluate each model on ImageNot.
 We observe that the final ranking is the same as on ImageNet.
-Moreover, the improvement that each model makes relative to prior models is roughly the same on ImageNot as it is on ImageNet.
+Moreover, each model’s improvement relative to prior models on ImageNot is roughly the same as on ImageNet.
 Figure 1 summarizes these results.
 The remarkable consistency of relative improvements suggests an intriguing counterfactual.
-Had ImageNot been the benchmark at the time of a new image classification model development, it could have induced the same improvements as ImageNet.
+Had ImageNot, or other large-scale datasets with different properties, been the benchmark at the time of a new image classification model development, it could have induced the same improvements as ImageNet.
+Crucially, we train from scratch on ImageNot and evaluate on an ImageNot test set, so we avoid the confound of model training that is pervasive in previous work to truly test external validity.
 A major hurdle in executing this comparison is that architectures developed for ImageNet do not readily train well on ImageNot.
@@
 ImageNot pretraining and transfer learning.
-ImageNet has proved useful for pretraining and transfer learning (Kornblith et al., 2019).
+ImageNet has proved useful for pretraining and transfer learning (kornblith2019better).
 We wanted to know if the same is true for ImageNot.
@@
   1. 1.
-    Pretrain on ImageN[o/e]t, finetune all layers on CIFAR-10, evaluate on CIFAR-10.
+    Pretrain on ImageNot/ImageNet, finetune all layers on CIFAR-10, evaluate on CIFAR-10.
   2. 2.
-    Pretrain on ImageN[o/e]t, finetune last layer on CIFAR-10, evaluate on CIFAR-10.
-We find that a model’s test accuracy on ImageN[o/e]t is a good predictor of fine-tunability and transferability.
-These experiments suggest that ImageNot shares with ImageNet its utility for pretraining and transfer learning.
+    Pretrain on ImageNot/ImageNet, finetune last layer on CIFAR-10, evaluate on CIFAR-10 – transfer.
+Similar to ImageNet, we find that a model’s test accuracy on ImageNot is also a good predictor of fine-tunability and transferability.
 To summarize, rankings of image classification models are surprisingly stable across significantly different datasets.
@@
 Even significant deliberate deviations from a dataset can maintain essential characteristics of model training and comparison.
-In particular, our findings suggest that the external validity of machine learning is, in important ways, significantly greater than previously assumed.
-Code.
+In particular, our findings suggest that the external validity of machine learning is, in important ways, significantly greater than previously assumed—Code.
 2 Discussion of related work
-The external validity of benchmark results for image classification has long been of concern.
-Indeed, benchmarks often contain biases that do not generalize to the data observed in the real world.
-Ponce et al.
-(2006) provide evidence of the narrow range of variation within existing benchmarks in the early stages of image classification research.
-Subsequently, Torralba and Efros (2011) compare multiple well-known image classification datasets, including SUN09 (Xiao et al., 2010), PASCAL Visual Object Classes (VOC) (Everingham et al., 2010), and Caltech-101 (Fei-Fei et al., 2004).
+Early work on dataset bias.
+The validity of benchmark results for image classification has long been of concern.
+Indeed, benchmarks often contain biases that do not generalize to data observed in the real world.
+ponce2006dataset provide evidence of the narrow range of variation within existing benchmarks in the early stages of image classification research.
+Subsequently, torralba2011unbiased compare multiple well-known image classification datasets, including SUN09 (xiao2010sun), PASCAL Visual Object Classes (VOC) (everingham2010pascal), and Caltech-101 (fei2004learning).
 They find evidence of selection bias, capture bias, and negative set bias in these datasets.
-The desire for a more representative benchmark for real-world image classification tasks gave birth to ImageNet, a large-scale image database (Deng et al., 2009).
-The extensive scale and relatively meticulous curation of ImageNet were thought to alleviate some of the limitations identified in earlier datasets, launching over a decade of active development in image classification, largely centered around improving accuracy on the ImageNet dataset.
-Still, there are concerns about the external validity of ImageNet results.
-Tsipras et al.
-(2020) studied the design choices of ImageNet and their effect on alignment between the ImageNet dataset and the broader underlying object recognition task.
+hardtrecht2022patterns Chapter 8 and (liberman2015reproducible) give background and history of datasets and benchmarks.
+The ImageNet era.
+The desire for a more representative benchmark for real-world image classification tasks motivated the creation of ImageNet, a large-scale image database (deng2009imagenet).
+The extensive scale and careful curation of ImageNet aimed to alleviate some of the limitations identified in earlier datasets.
+ImageNet was at the center of over a decade of active development in image classification, largely centered around improving accuracy on the ImageNet ILSVRC 2012 competition dataset.
+The observed progress on image classification can be well summarized by a set of models that span the active development of model architectures around ImageNet and a couple of modern models developed after the challenged ended in 2017: AlexNet (krizhevsky2012imagenet), VGG (simonyan2014very), Inception V3 (szegedy2015going), ResNet (he2016deep), DenseNet (huang2017densely), MobileNet V3 Large (howard2019searching), EfficientNet (tan2019efficientnet), ViT-B-16 (dosovitskiy2020image), and ConvNeXt (liu2022convnet).
+While other models were also developed during the so-called ImageNet era, we focus on models that were most impactful and represented a notable shift in architecture design—we include a discussion in Appendix A.
+Specifically, we considered the availability of model implementations across hubs such as TorchHub (pytorch_vision_classification), TensorFlow (abadi2015tensorflow), and Huggingface (wolf2020transformers) to be an indicator of popularity and, therefore, impact.
+Investigations about ImageNet.
+Not least due to its excessive use over the years, researchers registered concerns about the external validity of ImageNet results.
+pmlr-v119-tsipras20a studied the design choices of ImageNet and their effect on alignment between the ImageNet dataset and the broader underlying object recognition task.
 They concluded that ImageNet contains some inherent biases that cause misalignment.
@@
 However, they raised the issue that these biases make it difficult to identify actual improvements in object recognition.
-Xiao et al.
-(2020) identify one of these biases in that models tend to rely on the background of ImageNet images, which are correlated with the corresponding labels.
-Recht et al.
-(2019) investigated whether ImageNet classifiers generalize to a newly curated ImageNet test set—recall that by ImageNet, we generally refer to the ILSVRC 2012 ImageNet challenge.
+Xiao2020NoiseOS identify one of these biases in that models tend to rely on the background of ImageNet images, which are correlated with the corresponding labels.
+recht2019imagenet investigated whether ImageNet classifiers generalize to a newly curated ImageNet test set—recall that by ImageNet, we generally refer to the ILSVRC 2012 ImageNet challenge.
 They found that accuracies decrease strongly from the original the new test set, in the range of 11%-14%.
-Despite these findings, Recht et al.
-(2019)’s results showed that model rankings did not significantly change between the two ImageNet test sets, meaning, if model a performed better than model b on the original ImageNet test set, it also performed better on the new ImageNet test set.
-This speaks to the stability of model rankings under a slight distribution shift (new versus old test set) but leaves the question of how stable model rankings are under more drastic changes to the dataset.
-Kornblith et al.
-(2019) showed that for a set of popular image classification architectures, improvements on ImageNet are often a good predictor of transfer learning performance on other image classification datasets.
-In addition, beyond image classification, Huang et al.
-(2017b) showed that improvements in image classification on ImageNet often also predicted improvements in other computer vision tasks, such as image segmentation.
-While considerable research on the generalizability of ImageNet models has been centered around changes in accuracy under distribution shift, this work proposes a different perspective.
-We argue that for benchmarking purposes, providing robust model rankings is a more critical property of an effective benchmark than preserving accuracy or similar metrics.
-This, however, does not discount the importance of metrics such as accuracy on a well-thought-out and well-curated dataset.
-In Mark Liberman’s seminal 2015 lecture on reproducible research and the common task method, he described the role of benchmarks as an aligning mechanism for measuring progress (Liberman, 2015).
-We distinguish between two related functions of benchmarks in measuring progress: (i) Ranking, i.e., measuring relative progress, (ii) quantifying, i.e., absolute metrics such as accuracy in image classification.
-Although metrics can be useful, we argue they can be problematic targets for gauging external validity.
-Consider two identical benchmarks where we inserted a different fraction of random labels in each benchmark.
-The accuracy numbers may, therefore, be completely different, while the two benchmarks fundamentally measure the same thing.
-We, therefore, focus on ranking as the primary gauge of external validity.
-The observed progress on image classification can be well summarized by a set of models that span the active development of model architectures around ImageNet: AlexNet (Krizhevsky et al., 2012), VGG (Simonyan and Zisserman, 2014), ResNet (He et al., 2016), DenseNet (Huang et al., 2017a), EfficientNet (Tan and Le, 2019), and ConvNeXt (Liu et al., 2022).
-While other models were also developed during the so-called ImageNet era, we focus on models that were most impactful and represented a notable shift in architecture design – we include a discussion in Appendix A.
-Specifically, we considered the availability of model implementations across hubs such as TorchHub (PyTorch Team, 2024), TensorFlow (Abadi et al., 2015), and Huggingface (Wolf et al., 2020) to be an indicator of popularity and, therefore, impact.
-We note that while newer architectures such as Vision Transformers  (Devlin et al., 2019; Dosovitskiy et al., 2020) are now commonplace in the literature, we only consider architectures developed to improve ImageNet accuracy in our study.
-Many variants of ImageNet have been created to test the robustness of ImageNet models.
-Hendrycks and Dietterich (2019) introduced ImageNet-C/P, ImageNet with common corruptions.
-Hendrycks et al.
-(2020) introduced ImageNet-R, stylized distribution shifts on ImageNet.
-Li et al.
-(2022) introduced ImageNet-W, ImageNet with watermarks, a learning shortcut.
-Shirali and Hardt (2023) introduced LAIONet, an intended recreation of ImageNet from the LAION-400M dataset, in order to identify key differences between ImageNet and LAIONet.
-Our aim is distinctly different from these works.
-Rather than studying variants of ImageNet, we create an entirely different dataset.
-We aim to evaluate if this completely different dataset suggests the same model development trajectory, i.e., incremental rankings, as ImageNet.
-Accordingly, we introduce a new dataset, ImageNot, with entirely different classes and a unique curation process, drawing inspiration from LAIONet’s data-generating process (Shirali and Hardt, 2023).
-A distinct difference between our work and these previous works is that we train and test on this new dataset.
-This contrasts to (Recht et al., 2019; Shirali and Hardt, 2023), who compare test accuracies of the same ImageNet trained models on two different test sets with ImageNet classes.
-Our results compare the test accuracy of ImageNet-trained models on ImageNet’s test set with the test accuracy of ImageNot-trained models on ImageNot.
-This removes the confound of ImageNet training and the ImageNet classes, better isolating the external validity of model architectures independent of their training data.
-Additionally, we developed ImageNot rather than other datasets to ensure the strength of its contrast with ImageNet.
+barbu2019objectnet introduce a test set (ObjectNet) with real-world viewpoint, background, and rotation variation to assess model robustness and show that standard ImageNet-trained models suffer a 40–45% drop in top-1 accuracy on ObjectNet.
+Still, recht2019imagenet found that model rankings did not significantly change between the two ImageNet test sets: if model a performed better than model b on the original ImageNet test set, it generally also performed better on the new ImageNet test set.
+This speaks to the stability of model rankings under a slight distribution shift (new versus old test set), a form of internal validity.
+In other words, model rankings replicate under similar testing conditions.
+This form of internal validity is also supported by studies involving other variants of ImageNet.
+Hendrycks2019BenchmarkingNN introduced ImageNet-C/P, ImageNet with common corruptions.
+Hendrycks2020TheMF introduced ImageNet-R, stylized distribution shifts on ImageNet.
+Li2022AWD introduced ImageNet-W, ImageNet with watermarks, a learning shortcut.
+shirali2023makes introduced LAIONet, an intended recreation of ImageNet from the LAION-400M dataset, in order to identify key differences between ImageNet and LAIONet.
+External validity of ImageNet model rankings.
+From prior work, we know that model rankings satisfy internal validity, i.e., they replicate in similar conditions, whereas accuracy numbers do not.
+In this work, we ask a daring question about the external validity of model rankings: Do model rankings replicate in radically different conditions?
+Rather than studying variants of ImageNet and other smaller and similar datasets to ImageNet, we purposefully create a deliberately different dataset, in a sense, an anti-replication of ImageNet.
+Surprisingly, we find that models from the ImageNet era achieve the same relative improvements on ImageNot over prior models as they did on ImageNet.
+In contrast, we train models from scratch on the new dataset rather than reevaluate ImageNet-trained models on new test sets.
+One exception, perhaps closest to our study, is the important work of kornblith2019better, who examined how well the performance on ImageNet predicts performance on other vision datasets.
+They also train from scratch and find that on smaller and older datasets, such as DTD, VOC2007, or Caltech-101, model rankings do not replicate cleanly; however, they also find a strong correlation between ImageNet accuracy and accuracy on these datasets.
+While kornblith2019better considered standard computer vision datasets at a time, we purposefully create a dataset as different as possible from ImageNet from a data source (LAION) not available then.
+The key takeaway from  kornblith2019better could have been that ImageNet is special insofar as models developed on ImageNet transfer well.
+Our main conclusion, in contrast, is that from the perspective of model benchmarking, there may be nothing all that special about ImageNet other than its scale and diversity.
+In particular, from the benchmarking perspective, our study suggests that we do not even seem to need clean labels (beyer2020we).
+This surprising takeaway was confirmed in a theoretical study (dorner2024dont).
 3 Constructing ImageNot
-Figure 2: ImageNot caption x synset-gloss RoBERTa embeddings similarity for each class. We show the minimum, mean, median, and maximum similarity distribution across the 1000 classes.
+Figure 3: Overview of the ImageNot dataset curation pipeline.
+Starting from the LAION-5B corpus, the process selects all candidate noun synsets, removes classes overlapping with ImageNet or containing too few samples, filters out NSFW or ambiguous data, and ranks remaining classes by semantic (RoBERTa) and visual (CLIP) similarity.
+The final stage yields 1,000 ImageNot classes matched in scale but disjoint in concept from ImageNet, enabling evaluation of external validity.
+Figure 4: ImageNot caption x synset-gloss RoBERTa embeddings similarity for each class. We show some summary statistics of the similarity distribution across the 1000 classes.
 Figure 5: ImageNot caption x synset-gloss RoBERTa embeddings similarity distribution for each class. Each distribution corresponds to similarities for each class.
-We develop ImageNot, a dataset of the same scale as ImageNet’s ILSVRC-2012 – 1000 classes with 1000 examples per class – but without key design choices that have been thought critical in ImageNet’s effectiveness (Russakovsky et al., 2015).
+We develop ImageNot, a dataset of the same scale as ImageNet’s ILSVRC-2012—1000 classes with 1000 examples per class—but without key design choices that have been thought critical in ImageNet’s effectiveness (ILSVRC15).
 First, ImageNot has deliberately arbitrary classes constrained only to be strictly different from those in ImageNet.
@@
 There were no humans in the loop in selecting specific image and label pairs.
-We create the ImageNot dataset from the LAION-5B dataset, an academic, public massive scale collection of 5 billion image-text pairs crawled from the web and filtered by the CLIP model (Schuhmann et al., 2022; Radford et al., 2021).
-This is in contrast with the ILSVRC 2012 challenge, a subset of the ImageNet database which was constructed from Flickr image search results (Deng et al., 2009).
-Similar to ImageNet, ImageNot classes correspond to nouns in WordNet synsets found in LAION-5B captions, where WordNet is a large lexical database of English (Fellbaum, 1998; Miller, 1995; Hardeniya et al., 2016).
+We create the ImageNot dataset from the LAION-5B dataset, an academic, public massive scale collection of 5 billion image-text pairs crawled from the web and filtered by the CLIP model (schuhmann2022laion; radford2021learning).
+This contrasts the ILSVRC 2012 challenge, a subset of the ImageNet database which was constructed from Flickr search results (deng2009imagenet).
+ImageNot classes correspond to nouns in WordNet synsets found in LAION-5B captions, where WordNet is a large lexical database of English (fellbaum1998WordNet; miller1995WordNet; hardeniya2016natural).
 The following definitions are important for describing WordNet.
@@
 A WordNet tree represents a semantic hierarchy of synsets via an ‘IS-A’ structure, i.e., a synset is a ‘type’ of its synset parent, e.g., mammal $\rightarrow$ placental $\rightarrow$ carnivore $\rightarrow$ dog $\rightarrow$ working dog $\rightarrow$ husky.
-Each class in ImageNot represents a unique noun synset.
 At a high level, for each class, we select images from LAION-5B whose captions are similar to the synset corresponding to the class.
-We utilize RoBERTa (Liu et al., 2019) embeddings, an optimized version of BERT (Devlin et al., 2019), to estimate similarity between text sequences.
+We utilize RoBERTa (liu2019roberta) embeddings, an optimized version of BERT (devlin-etal-2019-bert), to estimate text sequence similarities.
 Since RoBERTa was never trained on ImageNet, this selection mechanism mitigates the concern that our image selection process may be biased in favor of ImageNet-like images.
-We note that creators of LAION-5B’s only included image and caption pairs that are sufficiently similar based on CLIP embeddings (Contrastive Language–Image Pre-training (Radford et al., 2021)), a multimodal deep learning model (Schuhmann et al., 2022).
+We note that creators of LAION-5B’s only included image and caption pairs that are sufficiently similar based on CLIP embeddings (Contrastive Language–Image Pre-training (radford2021learning)), a multimodal deep learning model (schuhmann2022laion).
 Constructing ImageNot.
-To choose the 1000 synsets (classes) that are included in ImageNot, we do the following (employing similar mechanics as Shirali and Hardt (2023)):
+In creating ImageNot, we build on the data-generating process developed by shirali2023makes.
+To choose the 1000 synsets (classes) that are included in ImageNot, available at Code, we take the following steps:
   1. 1.
@@
   5. 5.
-    Filter out synsets considered ‘Not Safe for Work’ (NSFW) by Song et al. (2021)’s NSFW classifier.
+    Filter out synsets considered ‘Not Safe for Work’ (NSFW) by song2021large’s NSFW classifier.
   6. 6.
@@
   7. 7.
-    Filter out synsets with an average top n caption x synset-gloss pair RoBERTa embeddings similarity (normalized inner product) less than 0.82  (Liu et al., 2019).
-We found 0.82 to be the highest reasonable choice to allow for sufficient examples in enough classes (Figure 3).
+    Filter out synsets with an average top n caption x synset-gloss pair RoBERTa embeddings similarity (normalized inner product) less than 0.82  (liu2019roberta).
+We found 0.82 to be the highest reasonable choice for sufficient examples in enough classes (Figure 5).
   8. 8.
-    Finally, select the top 1000 classes ordered by the average similarity between caption x synset-gloss pairs (Liu et al., 2019) (desideratum 4).
-Finally, we obtain ImageNot, a large-scale dataset that contrasts ImageNet.
-ImageNot has an average caption-label definition similarity of 0.91.
-Next, we compare ImageNot to ImageNet.
-ImageNot classes can be found at https://github.com/olawalesalaudeen/imagenot.
+    Finally, select the top 1000 classes ordered by the average similarity between the top 1100 class x image pairs (radford2021learning).
+Finally, we obtain ImageNot, a large-scale dataset that contrasts ImageNet. ImageNot has an average caption-label definition similarity of 0.91. Next, we compare ImageNot to ImageNet.
 3.1 ImageNot vs. ImageNet
-Figure 4: Synset depth in WordNet tree.
-Figure 5: Class Diversity. Average intra-class image similarity – inner product between normalized CLIP embeddings.
 Figure 6: Image Variance. Average inter-class image similarity – inner product between normalized CLIP embeddings.
-In this section, we delve into the differences between ImageNot and ImageNet, focusing on their design aspects.
-The primary distinctions lie in two areas: (i) the classes in ImageNot are arbitrarily different from those in ImageNet, and (ii) the attribution of a synset to an image in ImageNot is conducted without human intervention.
+Figure 7: ConvNext distinguishes ImageNot vs. ImageNet images with high accuracy. Error bars are over 3 trials.
+In this section, we examine the differences between ImageNot and ImageNet, focusing on their design aspects.
+The primary distinctions are: (i) the classes in ImageNot are deliberately far and different from those in ImageNet, and (ii) the attribution of a synset to an image in ImageNot is conducted without human intervention.
 We begin by examining the key differences between the two datasets.
@@
 In contrast, ImageNot is not designed to have densely populated subsets of the WordNet tree.
-Moreover, ImageNot synsets tend to be higher up in the WordNet tree (Figure 4) with ImageNot synsets having a median depth of 8 while ImageNet synsets have a median depth of 10 (more details about the WordNet subtrees can be found in Appendix B.1 Table 3).
+Moreover, ImageNot synsets tend to be higher up in the WordNet tree (Appendix B.1 Figure 11) with ImageNot synsets having a median depth of 8 while ImageNet synsets have a median depth of 10 (more details about the WordNet subtrees can be found in Appendix B Table 3).
 One consequence of this difference in median depth is that the average diversity of images in an ImageNot synset is higher than that of ImageNet.
@@
 The second significant not is the accuracy of image-synset pairs.
-In ImageNet, an average precision of 99.7% is achieved on a total of 80 synsets randomly sampled at different depths of the WordNet tree.
+In ImageNet, an average precision of 99.7% is achieved across 80 synsets randomly sampled at different depths of the WordNet tree (deng2009imagenet).
 Candidate images were collected from the internet by querying image search engines, and then human annotators were employed to verify each candidate image for a given synset, yielding relatively accurate image-synset pairs.
@@
 We rely entirely on similarities between embeddings from CLIP (image and caption similarities) and RoBERTa (caption and class definition similarities) to select class and image pairs.
-Consequently, ImageNot’s labels are far more noisy than ImageNet, which utilizes human annotators.
-This noisiness is noticeable in the accuracy observed when training on ImageNot – the best model achieves around 50%, which is drastically better than chance (0.1%), but lower than what the best model achieves on ImageNet (around 85%).
-We also see the effect of this process in the distribution of inter-class image similarities between ImageNot and ImageNet illustrated in Figure 6; the similarity in images corresponding to an ImageNet class tends to be higher than those in an ImageNot class.
-Figure 7: Number of synset siblings.
+Consequently, ImageNot’s labels are far noisier than ImageNet, which utilizes human annotators.
+We note that the ‘noise’ encompasses both incorrect labels due to automated curation and the greater natural variability of the images in ImageNot.
+This noisiness is noticeable in the accuracy observed when training on ImageNot – the best model achieves around 60%, which is drastically better than chance (0.1%) but lower than what the best model achieves on ImageNet (around 85%).
+We also see the effect of this process in the distribution of inter-class image similarities between ImageNot and ImageNet (Figure 7 and Appendix B.1 Figure 10); the similarity in images for an ImageNet class tends to be higher than those in an ImageNot class.
+Finally, we find that a classifier distinguishes ImageNet and ImageNot well—a common strategy to examine the extent to which two datasets differ (torralba2011unbiased; liu2024decade).
+We use a ConvNext (liu2022convnet), randomly initialized, as a binary classifier of the two datasets and achieve a 97.0% accuracy when trained on the full training sets.
+Even with only 1000 samples, accuracy is significantly above chance, and classification accuracy increases with more samples, as shown in Figure 7.
+Trials include new data resampling and model initialization.
+We sample training data uniformly for each class except for N/class=1000.
+The test set is sampled similarly; however, the test images are sampled from the ImageNot/ImageNet test sets.
 Similarities in ImageNot and ImageNet WordNet subtrees.
-Despite the differences in ImageNot, the subset of the WordNet trees spanned by their synsets are structurally similar.
-They share approximately the same number of nodes and edges, have similar depth, density and modularity, and share a 23% overlap in the parents of the datasets’ synsets – illustrated in Figure 7 (more details about the WordNet subtrees can be found in Appendix B.1 Table 3).
+Despite the differences in ImageNot/ImageNet, the subset of the WordNet trees spanned by their synsets is structurally similar.
+They share approximately the same number of nodes and edges, have similar depth, density and modularity, and share a 23% overlap in the parents of the datasets’ synsets – illustrated in Appendix B.1 Figure 12 (more on WordNet subtrees can be found in Appendix B.1 Table 3).
 4 ImageNot preserves model rankings
-(a) Models initialized with random weights. Train models on a dataset. Evaluate models on the same dataset’s held-out test set.
-(b) Relative progress of each model architecture. For a set of test accuracies $X$, relative progress is $X_{i}/X_{\text{AlexNet}}$.
-(c) Test accuracy compared to ImageNet test accuracy for each model architecture.
-Figure 8: ImageNet vs. ImageNot. We also show 95% confidence intervals for the regression estimates. Robust corresponds to average accuracy across many test-time distribution shifts.
 Our experimental design is centered around evaluating whether ImageNot maintains the model rankings established by ImageNet.
-Specifically, we want to know: if a new architectural advancement was evaluated on ImageNot instead of ImageNet, would it still be recognized as a significant development?
+Specifically, we want to know: if a new architectural advancement had been evaluated on ImageNot instead of ImageNet, would it still have been recognized as a significant development?
 Our focus is on a collection of model architectures that represent significant milestones in the evolution of ImageNet models.
-This includes AlexNet (Krizhevsky et al., 2012), VGG (Simonyan and Zisserman, 2014), ResNet (He et al., 2016), DenseNet (Huang et al., 2017a), EfficientNet (Tan and Le, 2019), and ConvNeXt (Liu et al., 2022).
-We motivate this set in section 2.
+This includes AlexNet (krizhevsky2012imagenet), VGG (simonyan2014very), Inception V3 (szegedy2015going), ResNet (he2016deep), DenseNet (huang2017densely), MobileNet (howard2019searching), EfficientNet (tan2019efficientnet), Vision Transformers (dosovitskiy2020image), and ConvNeXt (liu2022convnet)—see section 2 for motivating this set.
 For each model, we select the configuration that achieved the highest reported performance on ImageNet, such as ResNet152 over ResNet18.
@@
 We sort the test accuracy to establish a definitive order of performance in object recognition tasks, which we refer to as the ‘model ranking.’
-We also investigate model rankings on the CIFAR10 dataset, which consists of 60,000 32x32 colored images of 10 classes, with 5000 train images per class and 1000 test images per class (Krizhevsky et al., ; Krizhevsky et al., 2009).
+We also investigate ImageNet and ImageNot model performance for finetuning and transfer to the CIFAR10 dataset, which consists of 60,000 32x32 colored images of 10 classes, with 5000 training images per class and 1000 test images per class (cifar-10; krizhevsky2009learning).
 The learning task is 10-class image classification – Airplane, Automobile, Bird, Cat, Deer, Dog, Frog, Horse, Ship, and Truck.
-We note that ImageNet contains 50 test images per class (Deng et al., 2009) while ImageNot contains 100.
-We consider three types of rankings: (i) Standard.
-We initialize each model with random weight, then train each model on a dataset, and finally, we evaluate each model on the same dataset’s test set.
+We note that ImageNet contains 50 test images per class (deng2009imagenet) while ImageNot contains 100.
+We consider three types of experiments: (i) Random Initialization (From Scratch).
+We initialize each model with random weights, then train each model on a dataset, and finally, we evaluate each model on the same dataset’s test set.
 (ii) Fine-Tuning.
-We initialize each model with pre-trained on ImageN[o/e]t weights, then we fine-tune the model end-to-end (all weights updated) on CIFAR10, and finally, we evaluate it on CIFAR10’s test set.
+We initialize each model with pre-trained on ImageNot/ImageNet weights, then we fine-tune the model end-to-end (all weights updated) on CIFAR10, and finally, we evaluate it on CIFAR10’s test set.
 (iii) Transfer Learning.
-We initialize each model with pre-trained ImageN[o/e]t weights, then we update only the model’s last fully connected layer while training on CIFAR10, and finally, we evaluate it on CIFAR10’s test set.
+We initialize each model with pre-trained ImageNot/ImageNet weights, and then we update only the model’s last fully connected layer while training on CIFAR10.
+Finally, we evaluate it on CIFAR10’s test set.
 4.1 Model development
 ImageNet.
-Our reference rankings are derived from pre-trained models available on Torch Hub (Marcel and Rodriguez, 2010; Paszke et al., 2019).
+Our reference rankings are derived from pre-trained models available on Torch Hub (marcel2010torchvision; paszke2019pytorch).
 These models establish a ranking, which we consider our factual observations.
-In decreasing order of performance, the ranking is as follows: EfficientNet V2 Large, ConvNeXt Large, ResNet 152, DenseNet 161, VGG 19, ResNet 18, and AlexNet.
-For additional specifics regarding these pre-trained models, please refer to Appendix A.1.
-Training.
-We also utilize the implementation and pre-trained model weights from torch hub (Marcel and Rodriguez, 2010; Paszke et al., 2019), when applicable.
-For the first ranking category (no pretraining), we initialize each model with randomized weights and train end-to-end, meaning all model weights are updated with each backward pass.
+In decreasing order of performance, the ranking is as follows: EfficientNet V2 Large, ConvNeXt Large, Vision Transformer B-16, ResNet 152, Inception V3, DenseNet 161, MobileNet V3 Large, VGG 19, and AlexNet—Figure 1.
+Please refer to Appendix A.1 for additional specifics regarding these pre-trained models.
+We also utilize the implementation and pre-trained model weights from torch hub (marcel2010torchvision; paszke2019pytorch) for fine-tuning and transfer with ImageNet pretraining.
 On avoiding self-fulfilling prophecies.
-There is a subtle pitfall that one may fall into if not careful – the inadvertent confounding effect of knowing the desired ImageNet ranking.
+If not careful, one may fall into a subtle pitfall – the inadvertent confounding effect of knowing the desired ImageNet ranking.
 Specifically, an experimenter aware of the expected ranking might unconsciously adjust the process until a particular model is in the right place.
 This iterative tuning could falsely lead to a preservation of the known rankings.
-To mitigate this risk, we implemented several a priori restrictions on the experimenter’s approach for each ranking modality:
-  1. 1.
-    The experimenter can only tune the learning rate, learning rate scheduler hyper-parameters, and weight decay constant.
-  2. 2.
-    There is a shared hyper-parameter set across all models.
-  3. 3.
-    There is a fixed # of training runs that can be launched for the same maximum epoch.
-  4. 4.
-    The fixed hyper-parameter set can only be changed (expanded) if some model(s) do not train.
-See further discussion in Appendix A.2.
-This cautious approach introduces the primary limitation of our work.
-To avoid over-tuning and maintain objectivity, we refrain from employing extensive model tuning or leveraging all modern performance-enhancing techniques, such as data augmentation, label smoothing, etc.
-As a result, it is possible that we do not end up with the optimal training configuration for each model (Figure 8(c)).
+To mitigate this risk, we implemented a priori restrictions on the experimenter’s approach for each ranking modality (see Appendix A.1): (1) there is a fixed number of hyperparameter runs that can be launched per model (budget) and (2) the hyper-parameter search space is expanded for all models is expanded only if some model(s) do not train.
+For ImageNot, we start with the hyperparameters used to train reference ImageNet models to maintain objectivity.
+However, this means that our resulting training configuration for each model may not be optimal (Figure 8(a)-8(b)) relative to the decade-plus of tuning to obtain the best ImageNet models.
 4.2 Model Rankings
 ImageNot preserves model rankings and relative improvements.
-Figure 8(a) illustrates that the model rankings between ImageNot and ImageNet are remarkably consistent.
-In Figure 8(b), we compare each model’s performance improvement relative to AlexNet, the model that kick-started the deep learning era on ImageNet in 2012.
+Figure 1 illustrates that the model rankings between ImageNot and ImageNet are remarkably consistent.
+We compare each model’s performance improvement relative to AlexNet, the model that kick-started the deep learning era on ImageNet in 2012.
 This comparison reveals a notable parallel in the progression of accuracy improvements across both ImageNet and ImageNot.
-Not only do the ordinal rankings of models remain unchanged, the degree of performance enhancement is also consistent.
-In simpler terms, this suggests that if a new architecture like EfficientNet (Tan and Le, 2019) had been introduced in a different universe focused on ImageNot instead of ImageNet, we would have witnessed a comparable impressive advancement.
-The external validity of ImageNet improvements holds!
+Not only do the ordinal rankings of models remain unchanged, but the degree of performance enhancement is also consistent.
+In simpler terms, this suggests that if a new architecture like EfficientNet (tan2019efficientnet) had been introduced in a counterfactual universe focused on ImageNot instead of ImageNet, we would have witnessed a comparable impressive advancement.
+The external validity of ImageNet improvements follows.
 The ImageNet vs.
-ImageNot test accuracies have a linear fit with a slope of 0.46 and an adjusted $R^{2}$ of 0.98.
-ImageNet vs.
-CIFAR10 test accuracies have a linear fit with a slope of 0.18 and an adjusted $R^{2}$ of 0.93.
-We note that these slopes should not be interpreted the same way as those in Recht et al.
-(2019) because we compare test accuracies across datasets, i.e., different ranges in accuracy across models.
-(a) ImageN[o/e] pre-trained models. Train on CIFAR10. Evaluate on the CIFAR10. Similar to previous rankings, changes occur between architectures with overlapping error bars.
-(b) Fine-tuned on CIFAR10 test accuracy compared to standard ImageNet test accuracy for each model architecture.
-(c) Transfer Learning (fine-tune only last fully-connected layers) on CIFAR10 test accuracy compared to standard ImageNet test accuracy for each model architecture.
+ImageNot relative test accuracies have a linear fit with a slope of 1.01 and a Pearson R of 0.99.
+We note that these slopes should not be interpreted the same way as those in recht2019imagenet because we compare test accuracies of models with different training tasks/data with different difficulties.
+(a) Fine-tuned to CIFAR10—test accuracy compared to standard ImageNet test accuracy for each model architecture
+(b) Transfer Learning to CIFAR10—test accuracy compared to standard ImageNet test accuracy for each model architecture.
 Figure 8: ImageNet vs.
 ImageNot.
-Fine-tuning and Transfer Learning.
-Robust corresponds to average accuracy across many test-time distribution shifts.
-We also show 95% confidence intervals for the regression estimates.
-Figure 8(c) highlights the limitations of relying solely on the external validity of accuracy as the primary metric for the effectiveness of a benchmark.
-One may look at the drop in accuracy from ImageNet to ImageNot and conclude that ImageNet results have no external validity but this would be a rather narrow and inaccurate conclusion.
-From the standpoint of benchmarking, the more relevant observation and notion of external validity is that ImageNet accuracy is a strong predictor of ImageNot accuracy (Figure 8(c)), despite the significant difference between the two datasets.
-For each CIFAR10 ranking, we do three random restart restarts and compute the standard errors for each model’s test accuracy.
-Given that we opted to include architectures that achieve very similar accuracies on the ImageNet test set (such as EfficientNet-V2-L/ConvNeXt-L or ResNet152/DenseNet161), we expect potential training variance to affect the model, particularly given the size of CIFAR10.
-For instance, when training CIFAR10 from randomly initialized weights, EfficientNet-V2-L shows a slightly lower average test accuracy than ConvNeXt-L across the random restarts.
-However, their error bounds intersect, so the ranking is preserved up error margins.
+Fine-tuning and Transfer Learning with 95% confidence for regression estimates.
+Robust corresponds to average accuracy across test-time distribution shifts—error bars do not show for sufficiently small intervals.
+Our results also highlight the limitations of relying solely on the external validity of accuracy as the primary metric for the effectiveness of a benchmark.
+ImageNet model accuracies range from 0.57 (AlexNet) to 0.85 (EfficientNet V2 L), while much noisier ImageNot (Figure 2) model accuracies approximately range from 0.4 (AlexNet) to 0.6 (EfficientNet V2 L).
+One may look at the drop in accuracy from ImageNet to ImageNot and conclude that ImageNet results have no external validity, but this would be a rather narrow and inaccurate conclusion.
+From the standpoint of benchmarking, the more relevant observation and notion of external validity is that relative progression in ImageNet accuracy is nearly mirrored by progression in ImageNot accuracy, and model rankings are replicated exactly, despite the significant difference between the two datasets.
+For CIFAR10, we do three random restarts and compute the standard errors for each model’s test accuracy.
+We find that rankings are also preserved for CIFAR10 when training from scratch, up to error bars (Figure 9).
+kornblith2019better also evaluate CIFAR10 and find a similarly strong correlation between ImageNet accuracy and CIFAR10 accuracy.
+Given that we opted to include architectures that achieve very similar accuracies on the ImageNet test set (such as EfficientNet-V2-L/ConvNeXt-L or ResNet152/DenseNet161), we expect potential training variance to affect comparisons, particularly given the size of CIFAR10.
+For instance, when training CIFAR10 from randomly initialized weights or ImageNot (noisy), we observe overlapping error bars for a majority of the best models across random restarts.
+However, the consistency of relative rankings still holds on average.
 Our methodology restricts experimenters such that once a hyperparameter search space and the number of samples from the search space are fixed, they cannot be adjusted unless a model does not train.
@@
 However, we note that it also has some notable limitations, e.g., models with similar capabilities with overlapping error bars.
-ImageNot improvements yield improvements in fine-tuning and transfer learning.
-When fine-tuning pre-trained ImageN[o/e]t models on CIFAR10, we observe that model rankings are still preserved (Figure 9(a)).
-Within a fixed hyperparameter set, however, we find that some of the models, particularly those with the highest accuracies, have high variance.
-Nevertheless, both the upper bounds and lower bounds of the error margins increase as the accuracy on ImageNet increases (Figure 9(b)).
-Furthermore, despite the arbitrary nature of the ImageNot dataset, we observe that the representations it induces are still transferable – the accuracy of a model architecture on ImageNot is still a good predictor of the accuracy achieved by retraining the last layer of a model pre-trained on ImageNot on CIFAR10, (Figure 9(c)).
-However, we note that the ImageNot features are notably more noisy than ImageNet features when fine-tuning.
+ImageNot improvements predict improvements in fine-tuning and transfer learning.
+When fine-tuning pre-trained ImageNot/ImageNet models on CIFAR10, we observe that improvements in ImageNot predict improvements in finetunability of ImageNot models—similar to ImageNet and ImageNet models (Figure 8(a)).
+The ImageNot test accuracy and test accuracy of ImageNot models fine-tuned to CIFAR10 have a slope of 0.26 and a Pearson R of 0.83.
+Finetuned ImageNet has a slope and Pearson R of 0.29 and 0.95, respectively.
+Here, we find that, as a benchmark (our scope), ImageNot is counterfactually near-equivalent to ImageNet, but as a pretraining dataset, ImageNet is superior.
+When transferring ImageNot vs.
+ImageNet features, we find that ImageNet features are notably more performant (Figure 8(b)).
 This suggests that the noisiness of ImageNot does have some negative downstream effects on transfer learning.
-Nevertheless, we generally observe a positive trend in ImageNot test accuracy and transferability to CIFAR10.
-Rankings are preserved under robustness.
-We also find that model rankings remain consistent when we look at average performance across 17 distribution shifts applied to the ImageN[o/e]t test set, e.g., random blurring, cropping, and rotation.
-In this setting, our findings remain the same on average across these distribution shifts but with a consistent decrease in test accuracy for each model.
-This further shows that metrics can be a misleading definition of external validity as we observe changes in accuracy across different shifts.
-More details on rankings under robustness are in Appendix B.2.
+However, we still observe a positive trend in ImageNot test accuracy and transferability of ImageNot models to CIFAR10—slopes of 0.30 and 0.32 with ImageNot and ImageNet, respectively.
+Findings hold under robustness.
+We find that relative improvement also remains closely aligned between ImageNet and ImageNot when we look at average performance across 17 random data augmentations applied to the ImageNot/ImageNet test set (Figure 9; Appendix A.3), e.g., random blurring and cropping.
+ConvNeXt outperforms EfficientNet V2, however, this holds for both ImageNet and ImageNot, and MobileNet V3 performs marginally better than expected on ImageNot under these augmentations.
+Nevertheless, our results support the claim that preserving absolute metrics across datasets can be a misleading notion of external validity.
 Safety and ethical limitations of web-crawled datasets.
 While our results suggest that web-crawled datasets with minimal human intervention may be sufficient for benchmarking progress in architecture development, other ethical considerations are a concern.
-As illustrated by Thiel (2023), the internet contains many illegal images and images of illegal activity.
-Particularly, Thiel (2023) find evidence of Child Sexual Abuse Material (CSAM) in LAION-5B.
-We guard against this by using a NSFW language classifier (Song et al., 2021) to remove synsets that may be associated with unethical images on top of the default LAION-5B NSFW filter.
+As illustrated by thiel2023identifying, the internet contains many illegal images and images of illegal activity.
+Particularly, thiel2023identifying previously found evidence of Child Sexual Abuse Material (CSAM) in LAION-5B.
+We guard against this by using an NSFW language classifier (song2021large) to remove synsets that may be associated with unethical images on top of the default LAION-5B NSFW filter.
 We did not manually inspect every image in ImageNot, but we employed our model-based filtering and design choices to avoid problematic images as best as possible.
-Nevertheless, we do not release the images or URLs for ImageNot, but we provide ImageNot class names, a detailed description of our generation procedure, and our code so that our results can be reproduced.
+We provide ImageNot class names, a detailed description of our generation procedure, and our code so that our results can be easily reproduced.
 5 Conclusion
-Despite the significant differences in ImageNet and ImageNot, we find consistent preservation of model rankings across both datasets.
-Moreover, this consistency extends to the fine-tunability of pre-trained models and the transferability of pre-trained image embeddings.
-For benchmarking, we propose that providing robust model rankings should be a hallmark of a good benchmark.
-Furthermore, our results suggest that concerns about the external validity of model developments on ImageNet may be overstated.
+Despite the drastic differences in ImageNet and ImageNot, we find consistent preservation of model rankings across both datasets.
+Moreover, this consistency extends to pre-trained models’ fine-tunability and the transferability of pre-trained image embeddings.
+For benchmarking, we argue that providing robust model rankings, rather than absolute accuracy numbers, should be the primary function of a good benchmark.
+Consequently, our results suggest that concerns about the external validity of model developments on ImageNet as a benchmark may be overstated.
 Many domain experts attribute the rapid advancement in object recognition models to the ImageNet benchmark.
 From this work, two takeaways emerge to guide and motivate future benchmarks.
-First, a benchmark that produces robust model rankings, even if accuracy degrades under distribution shifts, should be considered effective.
-Second, the emphasis on clean labels and accuracy numbers may not be necessary to measure progress effectively, given that the noise is unbiased.
-The exact extent to which a benchmark needs to align with real-world image distributions to measure progress effectively remains a question for future work.
-Our findings indicate that this alignment might need to be less stringent than currently perceived in the machine learning community.
-Additional future work includes changing additional design choices of ImageN[o/e]t, e.g., moving away from wordnet noun synsets and using other attributes of or relationships between concepts in images.
-Another direction for future work is to examine the properties considered in this work for fine-tuning ImageNet and ImageNot-type models on other downstream tasks such as image segmentation.
+First, a benchmark that produces robust model rankings may be effective even if accuracy degrades under distribution shifts.
+Second, clean labels may not be necessary to measure progress effectively.
 Acknowledgements
@@
 Architecture Development Trajectory.
-AlexNet (Krizhevsky et al., 2012) demonstrated the efficacy of deep learning and many architectural advances such as convolutional filters, rectified linear units (ReLUs) as activation functions, and more.
+AlexNet (krizhevsky2012imagenet) demonstrated the efficacy of deep learning and many architectural advances such as convolutional filters, rectified linear units (ReLUs) as activation functions, and more.
 Additionally, its relative success on the ILSVRC 2012 is considered to have revitalized interest in neural networks, sparking a wave of research in deep learning.
-VGG (Simonyan and Zisserman, 2014) then implemented a simpler and deeper architecture with smaller convolutional filters, along with regularization techniques such as dropout, which yielded improved accuracy and generalization.
-ResNet (He et al., 2016) further improved upon the architecture of these deep learning models with the introduction of residual blocks, which allowed for the training of deeper networks while mitigating the vanishing gradient problem and achieving superior performance to its predecessor.
-DenseNet (Huang et al., 2017a) proposed to connect each layer to every other layer to maximize information flow between layers.
-EfficientNet (Tan and Le, 2019) then improved upon its predecessors with compound and systematic scaling of depth, width, and resolution, as well as combinations of novel optimization and learning strategies, without increasing computational cost.
-ConvNeXt (Liu et al., 2022) aimed to compete with the state-of-the-art performance of Vision Transformers (ViTs; (Dosovitskiy et al., 2020)) by combining components of preceeding ConvNet architectures, e.g., residual blocks and kernel sizes.
-A.1 Torch Hub Pretrained Models
-We use pretrained models from Torch Hub; we summarize their training details below (see PyTorch Team (2024) for additional details).
-Table 1: Default Hyperparameters
-Parameter Default Value
-batch size 32
-epochs 90
-lr 0.1
-momentum 0.9
-weight decay 1e-4
-lr step size 30
-lr gamma 0.1
-AlexNet, VGG19.
-Learning rate: 1e-2. Learning rate scheduler: StepLR.
-ResNet, DenseNet. Learning rate scheduler: StepLR.
-EfficientNet V2 L.
-Learning rate: 5e-1.
-Learning rate scheduler: Cosine annealing with linear warmup.
-Weight decay: 2e-5.
-EfficientNet V2 L also utilizes data augmentation, label smoothing, and other regularization techniques.
-ConvNeXt Large. Learning rate: 1e-3. Learning rate scheduler: Cosine annealing with linear warmup. ConvNeXt also utilizes data augmentation, label smoothing, and other regularization techniques.
-For both finetuning and transfer learning from ImageNet, we use these pretrained weights as our starting point.
-A.2 Trained Models
-We note that for ImageNot, the extent of our parameter tuning is significantly less than these pretrained ImageNet models (Table 2).
-Despite this, model rankings are still preserved.
-In general, when we train CIFAR10, we use the same set of hyperparameters, given that we can perform a broader search.
-This includes training from randomly initialized weights, finetuning, and transfer learning.
-For ImageNet, the tuning of these parameters we manual and not search-based – each new hyperparameter was added based on the loss curves of models that did not sufficiently train and domain expertise of which hyperparameters can be changed to allow for training.
-Table 2: Hyperparameters search space and Training Budgets for experiment. TL: Transfer Learning; FT: Finetuning; ST: Standard (train/test with randomly initialized weights).
+VGG (simonyan2014very) then implemented a simpler and deeper architecture with smaller convolutional filters, along with regularization techniques such as dropout, which yielded improved accuracy and generalization.
+Inception V3 szegedy2015going focused on multi-scale feature extraction by learning convolutions of different sizes in parallel.
+Additionally, auxiliary classifiers were added at intermediate layers to enhance gradient flow and act as implicit regularizers.
+ResNet (he2016deep) further improved upon AlexNet and VGG with the introduction of residual blocks, which allowed for the training of deeper networks while mitigating the vanishing gradient problem and achieving superior performance to its predecessor.
+DenseNet (huang2017densely) proposed to connect each layer to every other layer to maximize information flow between layers.
+MobileNet V3 Large focused on optimizing deep learning for mobile and embedded devices using depth-wise separable convolutions and squeeze-and-excitation (SE) modules (hu2018squeeze).
+EfficientNet (tan2019efficientnet) then improved upon its predecessors with compound and systematic scaling of depth, width, and resolution, as well as combinations of novel optimization and learning strategies, without increasing computational cost.
+Vision Transformers (ViT-B-16) apply a fundamentally different approach, replacing convolutional feature extractors with self-attention mechanisms; images are treated as sequences of patches, but self-attention across patches allows for long-range dependencies and flexible learning of spatial relationships.
+ConvNeXt (liu2022convnet) aimed to compete with the state-of-the-art performance of Vision Transformers (ViTs; (dosovitskiy2020image)) by combining components of preceding ConvNet architectures, e.g., residual blocks and kernel sizes.
+A.1 ImageNot Models
+For ImageNot, we utilize hyperparameters reported for the corresponding ImageNet model to match the model configuration.
+More details of the hyperparameters used are specified in our Code, with reference to https://github.com/pytorch/vision/tree/main/references/classification.
+We only change these parameters if models do not train.
+AlexNet.
+The AlexNet model uses a step scheduler and the SGD optimizer.
+The learning rate is set to 0.01 with a minimum learning rate of 0.0.
+The learning rate decay rate is 0.1, with no warmup epochs and an auxiliary learning rate decay after 30 epochs.
+The model is trained for 90 epochs with a momentum of 0.9 and a weight decay of 0.0001.
+The batch size is 32.
+ConvNeXt Large.
+The ConvNeXt Large model uses a cosine annealing with linear warmup scheduler and the AdamW optimizer.
+The learning rate is 0.001 with a minimum learning rate of 0.0.
+The learning rate decay rate is 0.1, and warmup is applied for 5 epochs.
+The model is trained for 600 epochs with a momentum of 0.9 and a weight decay of 0.05.
+The batch size is 128, and label smoothing is set to 0.1.
+Additionally, mixup is set to 0.2, cutmix to 1.0, repeated augmentation sampling (RA sampler) is enabled with 4 repetitions, and exponential moving average (EMA) decay is set to 0.99998 with 32 EMA steps.
+DenseNet-161.
+The DenseNet-161 model uses a step scheduler and the SGD optimizer.
+The learning rate is 0.1 with a minimum learning rate of 0.0.
+The learning rate decay rate is 0.01, with no warmup epochs and an auxiliary learning rate decay after 30 epochs.
+The model is trained for 90 epochs with a momentum of 0.9 and a weight decay of 0.0001.
+The batch size is 32.
+EfficientNet V2-L.
+The EfficientNet V2-L model uses a cosine annealing with linear warmup scheduler and the SGD optimizer.
+The learning rate is 0.5 with a minimum learning rate of 0.0.
+The learning rate decay rate is 0.1, and warmup is applied for 5 epochs.
+The model is trained for 600 epochs with a momentum of 0.9 and a weight decay of 0.00002.
+The batch size is 32, and label smoothing is set to 0.1.
+Additionally, mixup is set to 0.2, cutmix to 1.0, RA sampler is enabled with 4 repetitions, and EMA decay is set to 0.99998 with 32 EMA steps.
+Inception V3.
+The Inception V3 model uses a step scheduler and the SGD optimizer.
+The learning rate is 0.1 with a minimum learning rate of 0.0.
+The learning rate decay rate is 0.1, with no warmup epochs and an auxiliary learning rate decay after 30 epochs.
+The model is trained for 90 epochs with a momentum of 0.9 and a weight decay of 0.0001.
+The batch size is 32.
+Additionally, the auxiliary loss weight is set to 0.4.
+MobileNet V3-Large.
+The MobileNet V3-Large model uses a step scheduler and the RMSProp optimizer.
+The learning rate is 0.064 with a minimum learning rate of 0.0.
+The learning rate decay rate is 0.973, with no warmup epochs and an auxiliary learning rate decay after 2 epochs.
+The model is trained for 600 epochs with a momentum of 0.9 and a weight decay of 0.00001.
+The batch size is 128.
+ResNet-152.
+The ResNet-152 model uses a step scheduler and the SGD optimizer.
+The learning rate is 0.1 with a minimum learning rate of 0.0.
+The learning rate decay rate is 0.1, with no warmup epochs and an auxiliary learning rate decay after 30 epochs.
+The model is trained for 90 epochs with a momentum of 0.9 and a weight decay of 0.0001.
+The batch size is 32.
+ViT-B/16.
+The ViT-B/16 model uses a cosine annealing with a linear warmup scheduler and the AdamW optimizer.
+The learning rate is 0.001 with a minimum learning rate of 0.0.
+The learning rate decay rate is 0.033, and warmup is applied for 30 epochs.
+The model is trained for 300 epochs with a momentum of 0.9 and a weight decay of 0.3.
+The batch size is 512, and label smoothing is set to 0.11.
+Additionally, gradient clipping is set to 1.0, mixup is set to 0.2, cutmix to 1.0, RA sampler is enabled with 3 repetitions, and EMA decay is set to 0.99998 with 32 EMA steps.
+VGG-19.
+The VGG-19 model uses a step scheduler and the SGD optimizer.
+The learning rate is 0.01 with a minimum learning rate of 0.0.
+The learning rate decay rate is 0.1, with no warmup epochs and an auxiliary learning rate decay after 30 epochs.
+The model is trained for 90 epochs with a momentum of 0.9 and a weight decay of 0.0001.
+The batch size is 32.
+We utilize early stopping with patience of 10 based on held-out validation sample accuracy and set a maximum epoch of 1000 for all training.
+Additionally, to optimize training efficiency and reduce computational costs, we employ the Asynchronous Hyper-Band Scheduler from Ray Tune (li2020system; liaw2018tune), which enables us to further early-stop less promising hyperparameters.
+Compute Resource Estimate.
+We give a table of approximate times to reproduce our results below.
+Table 1: Compute Time. Note that this only accounts for the time it took to train the models used to produce the results in this paper.
+Dataset Results GPU-hrs GPUs
+ImageNot From Scratch 800 8 A100
+CIFAR10 From Scratch 40 2 A100
+CIFAR10/ImageNet Finetune 8 2 A100
+CIFAR10/ImageNot Finetune 12 2 A100
+CIFAR10/ImageNet Transfer Learning 4 2 A100
+CIFAR10/ImageNot Transfer Learning 6 2 A100
+A.2 CIFAR10 Models
+Table 2: Hyperparameters search space and Training Budgets for experiment.
+TL: Transfer Learning to CIFAR10; FT: Finetuning to CIFAR10; ST: Standard, randomly initialized weights trained only on CIFAR10.
 Dataset Hyperparameter Train Style Range Budget
@@
 CIFAR10 weight decay TL/FT/ST [1e-6, 5e-3] 24
-ImageNot batch size ST [256] 4
-ImageNot optimizer ST [SGD] 4
-ImageNot lr ST [5e-1, 1e-1, 1e-2, 1e-4] 4
-ImageNot lr decay ST [0.1] 4
-ImageNot lr scheduler ST [Linear Warm Start + Cosine Annealing] 4
-ImageNot momentum ST [0.9] 4
-ImageNot optimer warmup steps ST [5] 4
-ImageNot weight decay ST [2e-5] 4
-(*) For the reported batch size in Table 2, we use [128/32/32] for EfficientNet-V2-L and [256/64/64] for ConvNeXt-large.
-We utilize early stopping with a patience of 10 on the basis of held-out validation sample accuracy and set a maximum epoch of 1000 for all training.
-Additionally, to optimize training efficiency and reduce computational costs, we employ the Asynchronous Hyper-Band Scheduler from Ray Tune (Li et al., 2020; Liaw et al., 2018), which enables us to further early-stop less promising hyperparameters.
-Appendix B Additional Results and Discussion
+For both fine-tuning and transfer learning from ImageN[o/e]t, we use these pretrained weights as our starting point and then update weights either by fine-tuning or transfer learning.
+We also train end-to-end only on CIFAR10 with randomly initialized model weights.
+The hyperparameter details can be found in Table 2.
+A.3 Robustness
+We employ 17 distinct test-time data augmentations to evaluate the robustness of the developed models, all random except for the Identity augmentation.
+These augmentations are applied to the same set of held-out test examples.
+They include: (1) Identity, which uses the original image; (2) Gaussian Blur, applying a Gaussian blur to smooth the image and reduce noise; (3) Adjust Sharpness, for modifying image sharpness; (4) Affine, performing affine transformations such as scaling, rotations, and translations; (5) ColorJitter, adjusting brightness, contrast, saturation, and hue; (6) Contrast, altering contrast levels; (7) Equalize, for histogram adjustment and contrast balancing; (8) Grayscale, converting the image to grayscale; (9) Horizontal Flip, flipping the image horizontally; (10) Invert, inverting pixel values; (11) Perspective, applying perspective transformations; (12) Posterize, reducing the bit depth per color channel; (13) Resized Crop, cropping after resizing; (14) Rotation, rotating the image; (15) Sharpness, for sharpness adjustment; (16) Solarize, inverting colors post a threshold; and (17) Vertical Flip, flipping the image vertically.
+Appendix B Additional Results
+Figure 9: Final Model Ranking and Relative Improvement.
+Model rankings and relative improvements hold with models trained from scratch on the respective dataset, and evaluated on a held-out test set.
+For a set of test accuracies $X$, relative progress is $X_{i}/X_{\text{AlexNet}}$.
+ImageNet model accuracies range from 0.57 (AlexNet) to 0.85 (EfficientNet V2 L), while much noisier ImageNot model accuracies approximately range from 0.4 (AlexNet) to 0.6 (EfficientNet V2 L).
+CIFAR10 accuracies range from 0.89 (AlexNet) to 0.95 (ConvNeXt Large).
+The rankings for CIFAR10 are derived from means across three restarts; we note that ranking differences from ImageNet only occur for models with error bounds overlapping with the model in its ImageNet position.
+Thus, even for CIFAR10, the rankings are preserved up to error.
+Robust corresponds to average accuracy across many test-time distribution shifts.
+Figure 10: Class Diversity. Average intra-class image similarity – inner product between normalized CLIP embeddings.
 B.1 ImageNet vs. ImageNot WordNet subtree
+Table 3 summarizes the differences (or lack thereof) between properties of the ImageNet and ImageNot WordNet synsets (referenced in section 3.1).
+Figures 11 and 12 further depict the difference in ImageNot and ImageNet’s WordNet subtree structure.
+Figure 11: Synset depth in WordNet tree.
+Figure 12: Number of synset siblings.
 Table 3: ImageNot and ImageNet (ILSVRC 2012) WordNet tree statistics.
@@
 Max Depth 18 17 1.06
-Density 5.39e-4 5.60-4 0.96
+Density 5.39e-4 5.60e-4 0.96
 Modularity 0.94 0.93 1.01
@@
 Parent Overlap 23% 23%
-Table 3 summarizes the differences (or lack there of) between properties of the ImageNet and ImageNot WordNet synsets (referenced in section 3.1).
-B.2 Robustness
-Robustness Ranking.
-We employ 17 distinct test-time data augmentations to independently rank model accuracies, all random except for the Identity augmentation.
-These augmentations are applied to the same set of held-out test examples.
-They include: (1) Identity, which uses the original image; (2) Gaussian Blur, applying a Gaussian blur to smooth the image and reduce noise; (3) Adjust Sharpness, for modifying image sharpness; (4) Affine, performing affine transformations such as scaling, rotations, and translations; (5) ColorJitter, adjusting brightness, contrast, saturation, and hue; (6) Contrast, altering contrast levels; (7) Equalize, for histogram adjustment and contrast balancing; (8) Grayscale, converting the image to grayscale; (9) Horizontal Flip, flipping the image horizontally; (10) Invert, inverting pixel values; (11) Perspective, applying perspective transformations; (12) Posterize, reducing the bit depth per color channel; (13) Resized Crop, cropping after resizing; (14) Rotation, rotating the image; (15) Sharpness, for sharpness adjustment; (16) Solarize, inverting colors post a threshold; and (17) Vertical Flip, flipping the image vertically.
-We take the average across these augmentations to generate our robustness rankings. The model rankings are surprisingly stable on average across these 17 test-time augmentations (Figure 10-12).
-Figure 10: Models initialized with random weights. Train on ImageNet. Evaluate on ImageNet test set.
-We find that models trained on ImageNet maintain rankings on average when evaluated on the ImageNet test sets with test-time data corruption (Figure 10).
-Figure 11: Models initialized with random weights. Train on ImageNot. Evaluate on ImageNot test set.
-Additionally, we find that models trained on ImageNot maintain rankings on average when evaluated on the ImageNot test sets with test-time data corruption (Figure 11).
-However, across the corruptions, the rankings are less stable than what we observed for ImageNet.
-We hypothesize that this is a consequence of our guardrails on model tuning, which leads to sub-optimal models than those that could have been achieved with more extensive tuning.
-Figure 12: Models initialized with random weights. Train on CIFAR10. Evaluate on CIFAR10 test set.
-Finally, we find that models trained on CIFAR10 maintain rankings on average when evaluated on the CIFAR10 test sets with test-time data corruption (Figure 12).
-However, across the corruptions, the rankings are less stable than what we observed for ImageNet.
-We hypothesize that this is a consequence of our guardrails on model tuning, which leads to sub-optimal models than those that could have been achieved with more extensive tuning.